### v1

# Geometric foundations of thermodynamics in the quantum regime

Abstract
In this work, we present a comprehensive geometrical formulation of quantum thermodynamics based on contact geometry and principal fiber bundles. The quantum thermodynamic state space is modeled as a contact manifold, with equilibrium Gibbs states forming a Legendrian submanifold that encodes the fundamental thermodynamic relations. A principal fiber bundle over the manifold of density operators distinguishes the quantum state structure from thermodynamic labels: its fibers represent non-equilibrium configurations, and their unique intersections with the equilibrium submanifold enforce thermodynamic consistency. Quasistatic processes correspond to minimizing geodesics under the Bures–Wasserstein metric, leading to minimal dissipation, while the divergence of geodesic length toward rank-deficient states geometrically derives the third law. Non-equilibrium extensions, formulated through pseudo-Riemannian metrics and connections on the principal bundle, introduce curvature-induced holonomy that quantifies irreversibility in cyclic processes. In this framework, the thermodynamic laws in the quantum regime emerge naturally as geometric consequences.
I Introduction
Physics is geometry. Differential geometry [doCarmo1992Riemannian, ONeill1983SemiRiemannian, caja12_variedades, rudolph2012_1, rudolph2012_2] structures are fundamental to physical theories, from the symplectic geometry of classical mechanics [arnold1989mathematical] to the curved spacetime of general relativity [wald1984general] and the fiber bundle formulation of gauge theories [Bleecker1981Gauge, naber2011_1, naber2011_2, hamilton2017mathematical]. In thermodynamics [callen85_thermodynamics], contact geometry provides a natural and rigorous framework for equilibrium states and processes, as originally conceived by Gibbs and systematically developed in modern terms [bravetti15_contact, bravetti2017contact]. Here, the thermodynamic state space is a $(2n+1)$-dimensional contact manifold $(M,\eta)$, with the Gibbs 1-form $\eta$ encoding the first law, and equilibrium states forming a Legendrian submanifold $E$ where $\eta|_{E}=0$. Extending this geometric structure to the quantum regime [binder2018thermodynamics, strasberg2022quantum] presents significant conceptual and mathematical challenges. Recent works have explored quantum thermodynamics using “differential geometry” on the state space [scandi2019thermodynamic, anza2022geometric, tejero2025asymmetries, Bettmann2025Information]. However, there is not a proper geometrical formulation of the theory; just some processes analyzed with geometric perspective. A complete and unified geometric formulation of the full theory remains absent. Without a foundational geometric framework analogous to the classical case, analyses remain fragmented, and general principles are difficult to establish rigorously. Principal fiber bundles, central to gauge theories and topological phases in condensed matter [nakahara2018geometry, Bleecker1981Gauge], offer a powerful global structure for addressing these gaps. When applied to quantum thermodynamics, they furnish a principal bundle over the space of density operators, with fibers representing the thermodynamic coordinates associated with a fixed quantum state, thereby separating state evolution from thermodynamic variation and incorporating gauge-like symmetries.
In this paper, we synthesize these elements into a comprehensive geometrical framework extended to the quantum regime. The quantum thermodynamic state space is indeed a contact manifold $\mathcal{M}$, where thermal equilibrium states form the Legendrian submanifold $\mathcal{E}\subset\mathcal{M}$, satisfying the quantum first law. We introduce a principal fiber bundle over the space of density operators, mapping thermodynamic coordinates to states. Fibers enable analysis of relaxation paths, with equilibrium as unique intersections with the equilibrium submanifold. The Bures-Wasserstein metric on such submanifold [bengtsson2006geometry, oostrum22_bures] defines geodesics as quasistatic processes minimizing dissipation. Boundary effects in finite dimensions lead to geodesic incompleteness, recovering the third law geometrically. Non-equilibrium phenomena are addressed naturally in the theory, via pseudo-Riemannian extensions of the metric, compatible with $\ker\eta$, and the structure of the fiber bundle. The principal connection decomposes the tangent bundle, with curvature inducing holonomy in cyclic processes, a geometric source of irreversibility akin to gauge holonomies. This geometrization not only generalizes classical geometric thermodynamics but also provides mathematical rigor for quantum thermodynamics.
II Motivation: classical thermodynamics
Physics can be regarded as a collection of models that attempt to describe, always with some degree of idealization, specific aspects of the observable universe. Equilibrium thermodynamics exemplifies this as a robust physical theory centered on one of the most abstract idealizations: every system is assumed to be perpetually in thermodynamic equilibrium, implying that such a state would remain unchanged indefinitely. Paradoxically, this framework addresses phenomena like gas expansions, an out-of-equilibrium process. This is reconciled by positing that the evolution of the system proceeds exclusively through successive equilibrium states, necessitating an infinite duration for completion—a process termed quasistatic. Consequently, thermodynamic parameters vary smoothly, allowing well-defined derivatives with respect to an evolution parameter. These variations occur so gradually that they can be treated as infinitesimal, explaining why thermodynamic equations are typically expressed in differential forms. For instance, the first law is commonly written as $dU=TdS+\sum_{i}X_{i}dY_{i}+\sum_{j}\mu_{j}dN_{j}.$ The terms $dS$, $dY_{i}$, $dN_{j}$, for all $i,j$, can be interpreted as mere infinitesimals yielding finite integrals or, more insightfully, as differential 1-forms.
The latter perspective lends natural clarity to many thermodynamic concepts and results. Standard curricula on this topic usually distinguish state functions like entropy and internal energy—which depend solely on the system state when measured—from path-dependent quantities like heat or work. Thus, changes in internal energy or entropy are independent of the trajectory in thermodynamic parameter space, depending only on the endpoints, making them exact differentials. In contrast, integrals of heat or work generally vary with the path. A non-exact differential form may become exact when multiplied by an integrating factor, provided the form is closed. Assuming simply connected manifolds, closed and exact forms coincide. The prototypical example is heat: multiplying the infinitesimal amount $\delta Q$ by $1/T$ yields the exact differential $dS$.
Carathéodory’s theorem, often opaque in classical treatments, emerges straightforwardly geometrically: in any neighborhood of a thermodynamic state, certain states are inaccessible via quasistatic adiabatic processes. Geometrically, a process is a curve $\gamma:I\subseteq\mathbb{R}\to M$, where $M$ is the thermodynamic state space endowed with a differentiable manifold structure, as later explained. If $\delta Q\in\Lambda^{1}(M)$, the curve $\gamma$ is adiabatic when $\delta Q(\gamma^{\prime})=0$ everywhere. Since $dS=\delta Q/T$, it follows that $\delta Q=TdS$, so
$\delta Q(\gamma^{\prime})=TdS(\gamma^{\prime})=T\gamma^{\prime}(S)=T\frac{d}{dt}(S\circ\gamma)=0.$ (1)
Hence, quasistatic adiabatic processes cannot connect states of differing entropies.
The geometrization of classical thermodynamics provides a powerful framework for analyzing thermodynamic systems using tools from differential geometry. This approach represents thermodynamic systems as manifolds endowed with specific geometric structures, enabling a rigorous description of equilibrium states and thermodynamic processes. Here, we introduce the thermodynamic state space and its contact structure, establishing the foundation for extending these concepts to quantum thermodynamics in subsequent sections. The thermodynamic state space is formalized as a $(2n+1)$-dimensional manifold $M$, where $n$ is the number of independent extensive variables characterizing the system. This manifold is equipped with a contact structure defined by a 1-form $\eta$, satisfying the non-integrability condition $\eta\wedge(d\eta)^{n}\neq 0$, where $\wedge$ denotes the wedge (also exterior or Grassmann) product and $(d\eta)^{n}$ is the $n$-fold exterior product of $d\eta$. The contact structure ensures that $\eta$ defines a maximally non-integrable distribution that captures the constraints of thermodynamic processes. In physical terms, the thermodynamic state space is endowed with the Gibbs 1-form, typically expressed for a simple thermodynamic system as
$\eta=dU-TdS-\sum_{i=1}^{m}X_{i}dY_{i}-\sum_{j=1}^{k}\mu_{j}dN_{j},$ (2)
where $U$ is the internal energy, $S$ is the entropy, $Y_{i}$ are the extensive variables, $X_{i}$ are their conjugate intensive variables, $N_{j}$ are the particle numbers for $k$ species, and $T$ and $\mu_{j}$ are the temperature and chemical potentials, respectively, for $i=1,\ldots,m$, and $j=1,\ldots,k$. This 1-form encodes the first law of thermodynamics
$dU=TdS+\sum_{i=1}^{m}X_{i}dY_{i}+\sum_{j=1}^{k}\mu_{j}dN_{j},$ (3)
which relates changes in internal energy to work and heat exchanges in equilibrium processes. $(U,S,Y_{1},\ldots,Y_{m},N_{1},\ldots,N_{k},T,X_{1},\ldots,X_{m},\mu_{1},\ldots,\mu_{k})$ span the thermodynamic state space, with $\eta=0$ defining the equilibrium submanifold $E\subset M$, where thermodynamic states reside.
Following Callen [callen85_thermodynamics] and Bravetti et al. [bravetti15_contact], we formulate classical thermodynamics through the following postulates, which provide a mathematically consistent foundation for the geometric approach.
Postulate 1.
Equilibrium states of a thermodynamic system are fully characterized by a finite set of extensive variables: the internal energy $U$, generalized displacements $Y_{1},\ldots,Y_{m}$, and particle numbers $N_{1},\ldots,N_{k}$.
Postulate 2.
There exists a differentiable function
$\begin{split}S:E&\longrightarrow\mathbb{R}\\ (U,Y_{1},\ldots,Y_{m},N_{1},\ldots,N_{k})&\longmapsto S(U,Y_{1},\ldots,Y_{m},N_{1},\ldots,N_{k}),\end{split}$ (4)
called the entropy, defined on the space of equilibrium states $E$. This function is such that whenever an internal constraint of the system is removed, the new values of the extensive parameters are those which maximize the entropy over the manifold of constrained equilibrium states.
Postulate 3.
For a composite system comprising multiple subsystems, the total entropy is the sum of the entropies of the subsystems. Furthermore, the entropy $S$ is a differentiable, monotonically increasing function of the internal energy $U$.
These postulates have immediate geometric and physical implications. First, the equilibrium states form a finite-dimensional submanifold $E\subset M$, parameterized by the extensive variables $(U,Y_{1},\ldots,Y_{m},N_{1},\ldots,N_{k})$. On the equilibrium submanifold $E$, the Gibbs 1-form satisfies $\eta=0$, reflecting the first law in differential form. Finally, entropy is an intrinsic property for each particular subsystem and allows us to define the notion of temperature as the relation between the entropy and the energy, more concretely $T:=(\partial S/\partial U)^{-1}$. This relation follows from Postulate 2 and the monotonicity condition in Postulate 3.
This geometric framework provides a robust foundation for extending thermodynamic concepts to quantum systems. In classical thermodynamics, the thermodynamic state space and its contact structure capture the interplay between extensive and intensive variables. In quantum thermodynamics, we aim to generalize this structure to account for quantum states, described by density operators on a Hilbert space, and their thermodynamic properties. The equilibrium submanifold is replaced by a quantum state space, and the Gibbs 1-form will be adapted to incorporate quantum mechanical constraints.
The third law of thermodynamics, which governs the asymptotic behavior of entropy as temperature approaches absolute zero, is not considered in this geometric analysis in the classical regime. However, in this quantum case, it will arise naturally from the mathematical structure of the theory.
III Contact geometry
To establish the geometrical framework for quantum thermodynamics, we first review necessary concepts in contact geometry that underpin the thermodynamic state space and its extension to quantum systems. As stated, contact geometry provides a natural mathematical structure for describing thermodynamic equilibrium states and processes. This section introduces the essential definitions and results, setting the stage for their application to quantum thermodynamics in subsequent sections.
Definition 1.
Let $M$ be a smooth manifold of dimension $2n+1$. A 1-form $\eta\in\Lambda^{1}(M)$ is a contact form if it satisfies the non-integrability condition
$\eta\wedge(d\eta)^{n}\neq 0,$ (5)
where $(d\eta)^{n}$ is the $n$-fold exterior product of $d\eta$. This condition ensures that $\eta$ defines a volume form on $M$.
The non-integrability condition implies that contact forms exist only on odd-dimensional manifolds, as the wedge product $\eta\wedge(d\eta)^{n}$ is a $(2n+1)$-form, matching the dimension of $M$.
Proposition 1.
The contact form induces a contact distribution
$\mathscr{D}=\ker(\eta)=\{v\in TM\mid\eta(v)=0\},$ (6)
where $TM$ is the tangent bundle of $M$.
The distribution $\mathscr{D}$ has rank $2n$, and the condition $\eta\wedge(d\eta)^{n}\neq 0$ ensures that $\mathscr{D}$ is maximally non-integrable. It follows from Frobenius’ theorem that any submanifold of $M$ tangent to $\mathscr{D}$ has dimension at most $n$.
Definition 2.
A contact manifold is a pair $(M,\eta)$, where $M$ is a $(2n+1)$-dimensional smooth manifold and $\eta$ is a contact form on $M$.
In the context of thermodynamics, the thermodynamic state space is modeled as a contact manifold, with the Gibbs 1-form, Eq. (2), serving as the contact form. The contact structure encodes the thermodynamic relations, such as the first law.
Definition 3.
An $n$-dimensional submanifold $\mathcal{L}\subset M$ of a contact manifold $(M,\eta)$ is Legendrian if $\eta|_{\mathcal{L}}=0$.
In thermodynamics, Legendrian submanifolds correspond to equilibrium submanifolds $E\subset M$, where the Gibbs 1-form satisfies $\eta=0$. These submanifolds represent the set of equilibrium states, parameterized by variables such as internal energy $U$, entropy $S$, volume $V$, and particle numbers $N_{i}$.
Since $\eta\wedge(d\eta)^{n}\neq 0$, the differential $d\eta$ must have one-dimensional kernel, which is transverse to $\mathscr{D}$. This condition yields a unique vector field on the manifold, which governs the dynamics.
Definition 4.
The Reeb vector field $\mathcal{R}_{\eta}$ on a contact manifold $(M,\eta)$ is the unique vector field satisfying the conditions $\eta(\mathcal{R_{\eta}})=1$, and $d\eta(\mathcal{R_{\eta}},\cdot)=0$.
In thermodynamics, the Reeb field is associated with the evolution of thermodynamic processes. An essential result in contact geometry is the existence of canonical coordinates, which simplifies the local description of the contact structure.
Theorem 1 (Darboux).
Let $(M,\eta)$ be a $(2n+1)$-dimensional contact manifold. For any point $p\in M$, there exist local coordinates $(x_{1},\ldots,x_{n},y_{1},\ldots,y_{n},z)$ in a neighborhood of $p$ such that the contact form is expressed as
$\eta=dz-\sum_{i=1}^{n}y_{i}dx_{i}.$ (7)
In thermodynamics, Darboux theorem ensures that the thermodynamic state space can be locally described using coordinates that align the Gibbs 1-form with the contact structure. For example, for a simple system, coordinates can be chosen such that $\eta=dU-TdS+pdV$. This canonical form facilitates the analysis of thermodynamic relations and their extension to quantum systems.
IV Connection to quantum thermodynamics
In this section, we construct the quantum thermodynamic state space as a contact manifold, generalizing the classical thermodynamic state space to quantum systems. The quantum thermodynamic state space provides a geometric framework for describing quantum equilibrium states, encoded by Gibbs states, and their thermodynamic relations.
Definition 5.
Let $\mathcal{H}$ be an $m$-dimensional complex Hilbert space, so that $\mathcal{H}\cong\mathbb{C}^{m}$ for $m<\infty$, and let $\mathscr{B}(\mathcal{H})$ be the set of bounded operators on $\mathcal{H}$. A quantum state is any $\rho\in\mathscr{B}(\mathcal{H})$ such that $\rho$ is self-adjoint, positive semidefinite and has unit trace.
See [gustafson03_mathematical, takhtadzhian08_quantum] for further details. In finite dimensions, $\mathscr{B}(\mathcal{H})\cong M_{m}(\mathbb{C})$. By virtue of this isomorphism, we identify states with unit trace self-adjoint matrices and introduce the following topological and differentiable structure.
Proposition 2.
The set of quantum states $\mathcal{D}=\{\rho\in\mathscr{B}(\mathcal{H})\mid\rho\ \mathrm{quantum\ state}\}$ is a compact convex subset of the hyperplane $\Pi=\{\pi\in\mathscr{B}(\mathcal{H})\mid\pi\ \mathrm{self}{-}\mathrm{adjoint},\ \operatorname{\mathrm{tr}}(\pi)=1\}\cong\mathbb{R}^{m^{2}-1}$.
Proof.
Let $\rho_{1},\rho_{2}\in\mathcal{D}$ and $\lambda\in[0,1]$. Then $\lambda\rho_{1}+(1-\lambda)\rho_{2}$ is self-adjoint (as a convex combination of self-adjoint operators), positive semidefinite (since the positive semidefinite cone is convex: the eigenvalues of the combination lie in the convex hull of the eigenvalues of $\rho_{1}$ and $\rho_{2}$, hence remain non-negative), and satisfies $\operatorname{\mathrm{tr}}[\lambda\rho_{1}+(1-\lambda)\rho_{2}]=1$ by linearity of the trace. Thus, $\mathcal{D}$ is convex. ∎
Definition 6.
The relative interior of $\mathcal{D}$ is the open set of positive definite states, defined as $\mathcal{D}^{\circ}:=\operatorname{int}_{\Pi}(\mathcal{D})=\{\rho\in\Pi\mid\rho>0\}$.
Remark 1.
The positive definite states are those with full rank, i.e. $\mathrm{rank}(\rho)=m$, for all $\rho\in\mathcal{D}^{\circ}$.
Proposition 3.
$\mathcal{D}^{\circ}$ is a smooth manifold of dimension $m^{2}-1$.
Proof.
The manifold structure of $\mathcal{D}^{\circ}$ is trivial [caja12_variedades]. ∎
Definition 7.
The boundary of the state space is $\partial\mathcal{D}:=\mathcal{D}\setminus\mathcal{D}^{\circ}=\{\rho\in\mathcal{D}\mid\mathrm{rank}(\rho)<m\}$.
Remark 2.
The boundary $\partial\mathcal{D}$ contains all rank-deficient states, i.e. $\mathrm{rank}(\rho)<m$, including the pure states, i.e. states with rank $1$.
The subsequent analysis is restricted to the interior $\mathcal{D}^{\circ}$. Although boundary effects must be considered in general equilibrium manifold theory due to finite-dimensional constraints, they play no role in the present context, as will be shown in Sec. IX.
Definition 8.
Let $\mathcal{H}$ be a Hilbert space. An observable $A$ is a self-adjoint bounded operator on $\mathcal{H}$. For any quantum state $\rho\in\mathcal{D}$, the expectation value of $A$ in the state $\rho$ is defined as
$a=\langle A\rangle_{\rho}=\operatorname{\mathrm{tr}}(\rho A)\in\mathbb{R}.$ (8)
A finite collection $\{A_{1},\ldots,A_{n}\}$ of observables is called a set of observables.
For later convenience, we introduce a finite set of observables $\boldsymbol{A}=\{A_{i}\in\mathscr{B}(\mathcal{H})\mid A_{i}\text{ observable},i=1,\ldots,n\}$. We associate to this set a collection of real-valued conjugate variables $\boldsymbol{\lambda}=\{\lambda_{i}\in\mathbb{R}\mid i=1,\ldots,n\}$, acting as smooth coordinates on the thermodynamic state space. Mimicking classical thermodynamics, the interest resides in defining a contact manifold for the quantum states, for them to be labeled by several coordinates by Theorem 1.
Theorem 2 (Quantum thermodynamic state space).
The quantum thermodynamic state space is a smooth manifold $\mathcal{M}\cong\mathbb{R}^{2n+1}$, with global coordinates $(S,\boldsymbol{a},\boldsymbol{\lambda})=(S,a_{1},\ldots,a_{n},\lambda_{1},\ldots,\lambda_{n})$, endowed with a contact form
$\eta=dS-\sum_{i=1}^{n}\lambda_{i}da_{i}\in\Lambda^{1}(\mathcal{M}).$ (9)
Proof.
The diffeomorphism $\mathcal{M}\cong\mathbb{R}^{2n+1}$ is immediate from the coordinate chart. To verify that $\eta$ is a contact form, compute the exterior derivative
$d\eta=d\left(dS-\sum_{i=1}^{n}\lambda_{i}da_{i}\right)=-\sum_{i=1}^{n}d\lambda_{i}\wedge da_{i},$ (10)
since $d(dS)=0$ and $d(\lambda_{i}da_{i})=d\lambda_{i}\wedge da_{i}$. Then
$(d\eta)^{n}=(-1)^{n}\bigwedge_{i=1}^{n}\left(d\lambda_{i}\wedge da_{i}\right).$ (11)
and the wedge product is
$\eta\wedge(d\eta)^{n}=dS\wedge(-1)^{n}\bigwedge_{j=1}^{n}(d\lambda_{i}\wedge da_{i}).$ (12)
The set $\{dS,da_{1},\dots,da_{n},d\lambda_{1},\dots,d\lambda_{n}\}$ is a basis of $T^{*}\mathcal{M}$, so the right-hand side is a non-vanishing volume form. Hence $\eta\wedge(d\eta)^{n}\neq 0$ everywhere and $(\mathcal{M},\eta)$ is a contact manifold. ∎
To characterize quantum states in thermodynamic equilibrium, we introduce the following postulate, which specifies their mathematical form.
Postulate 4.
A quantum state $\rho\in\mathcal{D}$ is in thermodynamic equilibrium with respect to a set of observables $\{A_{1},\dots,A_{n}\}\subset\mathscr{B}(\mathcal{H})$ if it is given by
$\rho_{\boldsymbol{\lambda}}=\frac{\exp\left(-\sum_{i=1}^{n}\lambda_{i}A_{i}\right)}{Z(\boldsymbol{\lambda})},$ (13)
with
$Z(\boldsymbol{\lambda})=\operatorname{\mathrm{tr}}\left[\exp\left(-\sum_{i=1}^{n}\lambda_{i}A_{i}\right)\right],$ (14)
where $\boldsymbol{\lambda}=(\lambda_{1},\dots,\lambda_{n})\in\mathbb{R}^{n}$. Such states are known as Gibbs states, and maximize the entropy for a given set of observables $\{A_{1},\dots,A_{n}\}\subset\mathscr{B}(\mathcal{H})$.
Remark 3.
Since the exponential of a matrix is always invertible, it has maximum rank, meaning that all states of the form given by Eq. (13) lie in $\mathcal{D}^{\circ}$.
To establish a connection between the contact structure and the quantum viewpoint, we define a state function that maps thermodynamic coordinates to quantum states.
Definition 9.
Let $\mathcal{D}$ be the set of all density operators on a finite-dimensional Hilbert space $\mathcal{H}\cong\mathbb{C}^{m}$. The state function $\Xi:\mathcal{M}\to\mathcal{D}$ assigns to each thermodynamic point $(S,\boldsymbol{a},\boldsymbol{\lambda})$ the density operator
$\Xi(S,\boldsymbol{a},\boldsymbol{\lambda})=\frac{\exp\left(-\sum_{i=1}^{n}\mu_{i}(S,\boldsymbol{a},\boldsymbol{\lambda})A_{i}\right)}{\operatorname{\mathrm{tr}}\left[\exp\left(-\sum_{i=1}^{n}\mu_{i}(S,\boldsymbol{a},\boldsymbol{\lambda})A_{i}\right)\right]},$ (15)
where the smooth functions $\mu_{i}:\mathcal{M}\to\mathbb{R}$ satisfy the equilibrium consistency condition, i.e. given the embedding
$\begin{split}\iota:\mathbb{R}^{n}&\longrightarrow\mathcal{M}\\ \boldsymbol{\lambda}&\longmapsto\left(S(\boldsymbol{\lambda}),\boldsymbol{a}(\boldsymbol{\lambda}),\boldsymbol{\lambda}\right),\end{split}$ (16)
where
$\begin{split}a_{i}(\boldsymbol{\lambda})&=\langle A_{i}\rangle_{\rho_{\boldsymbol{\lambda}}}=-\frac{\partial\log Z}{\partial\lambda_{i}},\\ S(\boldsymbol{\lambda})&=-\operatorname{\mathrm{tr}}\left(\rho_{\boldsymbol{\lambda}}\log\rho_{\boldsymbol{\lambda}}\right)=\log Z(\boldsymbol{\lambda})+\sum_{i=1}^{n}\lambda_{i}a_{i}(\boldsymbol{\lambda}),\end{split}$ (17)
we have
$(\mu_{i}\circ\iota)(\boldsymbol{\lambda})=\lambda_{i},$ (18)
for all $\boldsymbol{\lambda}\in\mathbb{R}^{n}$, for $i=1,\ldots,n$.
The functions $\mu_{i}$ generalize the conjugate parameters $\lambda_{i}$ to non-equilibrium points, ensuring that $\Xi$ is a smooth map across $\mathcal{M}$. On the equilibrium submanifold, where $\mu_{i}=\lambda_{i}$, the state reduces to a Gibbs state, which maximizes the entropy for given expectation values $a_{i}=\operatorname{\mathrm{tr}}(A_{i}\rho)$.
Proposition 4.
The image of the previous embedding is an $n$-dimensional Legendrian submanifold $\mathcal{E}=\iota(\mathbb{R}^{n})\subset\mathcal{M}$.
Proof.
On $\mathcal{E}$ the functions $\mu_{i}$ reduce to $\lambda_{i}$ by Eq. (18), so
$\Xi\bigl(S(\boldsymbol{\lambda}),\boldsymbol{a}(\boldsymbol{\lambda}),\boldsymbol{\lambda}\bigr)=\rho_{\boldsymbol{\lambda}}.$ (19)
Differentiating the identity $S=\log Z+\sum_{i}\lambda_{i}a_{i}$ yields
$dS=\sum_{i}\lambda_{i}da_{i}+\sum_{i}a_{i}d\lambda_{i}+\sum_{i}\frac{\partial\log Z}{\partial\lambda_{i}}d\lambda_{i}.$ (20)
The partial derivative $\partial\log Z/\partial\lambda_{i}=-a_{i}$, hence the last two terms cancel and $dS=\sum_{i=1}^{n}\lambda_{i}da_{i}$. The pull-back of the contact form is therefore $\iota^{*}\eta=dS-\sum_{i}\lambda_{i}da_{i}=0$. Since $\dim\mathcal{E}=n$ and $\iota^{*}\eta=0$, $\mathcal{E}$ is Legendrian. ∎
In thermodynamic equilibrium, i.e. on $\mathcal{E}$, the coordinates have physical interpretations. The coordinate $S$ is the von Neumann entropy $S=-\operatorname{\mathrm{tr}}(\rho\log\rho)$. Each $a_{i}=\operatorname{\mathrm{tr}}(A_{i}\rho)$ is the expectation value of the observable $A_{i}$. The conjugate parameters $\lambda_{i}$ play the role of intensive variables, e.g. inverse temperatures or chemical potentials. The contact relation $dS=\sum_{i}\lambda_{i}da_{i}$ is the quantum version of the classical first law in the entropy representation.
Remark 4 (First law of quantum thermodynamics).
The quantum first law is encoded in the contact form $\eta=dS-\sum_{i=1}^{n}\lambda_{i}da_{i}$ on the quantum thermodynamic state space $\mathcal{M}$. On the Legendrian submanifold $\mathcal{E}$ of equilibrium states, $\eta|_{\mathcal{E}}=0$ implies
$dS=\sum_{i=1}^{n}\lambda_{i}da_{i},$ (21)
expressing the differential change of the entropy in terms of variations in expectation values of observables and intensive parameters. This generalizes the classical first law to quantum systems.
Remark 5.
Henceforth, thermodynamic entropy in the quantum regime is quantified by the von Neumann entropy, defined as $S(\rho)=-\operatorname{\mathrm{tr}}(\rho\ln\rho)$, for all $\rho\in\mathcal{D}^{\circ}$. Although the precise interpretation of quantum thermodynamic entropy remains a subject of active debate, for our purposes, the specific interpretation is immaterial, provided the entropy is a smooth, positive and concave function $S:\mathcal{D}^{\circ}\to\mathbb{R}$ on the manifold of density operators. These conditions ensure compatibility with the differential geometric framework developed herein.
The functions $\mu_{i}:\mathcal{M}\to\mathbb{R}$, $i=1,\dots,n$ allow $\Xi$ to be defined smoothly everywhere in $\mathcal{M}$, generalizing Gibbs states to non-equilibrium points. On the equilibrium submanifold $\mathcal{E}$, the consistency condition given by Eq. (18) enforces $\mu_{i}\left(S(\boldsymbol{\lambda}),\boldsymbol{a}(\boldsymbol{\lambda}),\boldsymbol{\lambda}\right)=\lambda_{i}$, for all $i=1,\ldots,n$ so that $\Xi\bigl(S(\boldsymbol{\lambda}),\boldsymbol{a}(\boldsymbol{\lambda}),\boldsymbol{\lambda}\bigr)=\rho_{\boldsymbol{\lambda}}$ recovers the standard Gibbs state as in Eq. (13). Off equilibrium, i.e. for points away from $\mathcal{E}$, the functions $\mu_{i}$ are not constrained by the Gibbs form and may be chosen flexibly to model non-equilibrium dynamics. For example, the correspondence $\mu_{i}=\lambda_{i}+f_{i}(S,\boldsymbol{a},\boldsymbol{\lambda})$, where each $f_{i}\in C^{\infty}(\mathcal{M})$ vanishes on $\mathcal{E}$, i.e. $f_{i}|_{\mathcal{E}}=0$, provides a valid extension, for all $i=1,\ldots,n$.
Remark 6.
A naive attempt to define $a_{i}=\operatorname{\mathrm{tr}}(A_{i}\Xi(S,\boldsymbol{a},\boldsymbol{\lambda}))$ globally on $\mathcal{M}$ would render $\Xi$ implicitly dependent on its own output, leading to circularity. This is avoided by treating the coordinates $(S,\boldsymbol{a},\boldsymbol{\lambda})$ as independent in $\mathcal{M}\cong\mathbb{R}^{2n+1}$. The physical constraint by $a_{i}=\operatorname{\mathrm{tr}}(A_{i}\rho)$ is imposed only on the equilibrium submanifold $\mathcal{E}$, where it is automatically satisfied by the Gibbs construction: see Proposition 4. Elsewhere, $\boldsymbol{a}$ represents target or virtual expectation values, not necessarily realized by $\Xi$, enabling the modeling of constrained or fictitious ensembles.
V Quantum thermodynamic fiber bundle
Under appropriate circumstances, the state function introduced in Definition 9 induces a rich geometric structure: $\mathcal{M}$ is a fiber bundle over the base space of quantum states, with fibers encoding all thermodynamic configurations compatible with a given density operator. This framework unifies equilibrium and non-equilibrium quantum thermodynamics within a single differential-geometric object.
Let us suppose that the manifold of Gibbs states generated by the fixed observables $\{A_{1},\dots,A_{n}\}$, which is given by $\mathcal{B}=\bigl\{\rho_{\boldsymbol{\lambda}}\big|\boldsymbol{\lambda}\in U\subseteq\mathbb{R}^{n}\bigr\}\subseteq\mathcal{D}^{\circ}$, has maximal dimension $n$. Then, the correspondence $\Xi$ is a smooth map onto the $n$-dimensional submanifold $\mathcal{B}$. Additionally, if $\Xi$ is a submersion, i.e. the differential $d\Xi_{p}:T_{p}\mathcal{M}\to T_{\Xi(p)}\mathcal{B}$ is surjective for all $p\in\mathcal{M}$ we can claim the following:
Definition 10.
Let $\mathcal{M}\cong\mathbb{R}^{2n+1}$ be the quantum thermodynamic state space with coordinates $(S,\boldsymbol{a},\boldsymbol{\lambda})$, and let $\mathcal{B}$ be the manifold of Gibbs states generated by the fixed observables $\{A_{1},\dots,A_{n}\}$. The quantum thermodynamic fiber over a Gibbs state $\sigma\in\mathcal{B}$ is the preimage
$F_{\sigma}:=\Xi^{-1}(\sigma)=\bigl\{(S,\boldsymbol{a},\boldsymbol{\lambda})\in\mathcal{M}\;\big|\;\Xi(S,\boldsymbol{a},\boldsymbol{\lambda})=\sigma\bigr\}.$ (22)
If $\Xi$ is a submersion, the fiber $F_{\sigma}$ is a smooth submanifold of $\mathcal{M}$ with dimension
$\dim F_{\sigma}=\dim\mathcal{M}-\dim\mathcal{B}=(2n+1)-n=n+1.$ (23)
Remark 7.
The conditions required for the previous definition to make sense, namely that the submanifold $\mathcal{B}$ have dimension $n$ and that the map $\Xi$ be a submersion, are not guaranteed to hold. A first immediate requirement for $\mathcal{B}$ to have maximal dimension is that the observables $\{A_{1},\dots,A_{n}\}$ be linearly independent. This ensures that the submanifold generated by the exponential of the linear combinations of such observables has the correct dimension $n$. However, dividing by the trace may entail a reduction in the dimensionality of the submanifold. For example, if the set of observables only has one element $\{A_{1}\}$ and $A_{1}$ is a multiple of the identity matrix, the set $\mathcal{B}$ contains just one element. On the contrary, if $A_{1}$ is a $n\times n$ matrix having $n$ distinct eigenvalues, $\mathcal{B}$ has dimension $n$ and $\Xi$ is a submersion. Cases where the set of observables has more than one element have to be studied separately.
Definition 11.
A quantum thermodynamic fiber bundle is a tuple $(\mathcal{M},\mathcal{B},\Xi,F)$, where $\mathcal{M}$ is the quantum thermodynamic state space (total space), with $\dim=2n+1$; $\mathcal{B}\subseteq\mathcal{D}^{\circ}$ is the base space, with $\dim=n$; $F\cong\mathbb{R}^{n+1}$ is the fiber; and $\Xi:\mathcal{M}\to\mathcal{B}$ is the smooth projection (state function). The bundle satisfies local triviality: for every $\sigma\in\mathcal{B}$, there exists a neighborhood $U\subset\mathcal{B}$ and a diffeomorphism $\phi:\Xi^{-1}(U)\to U\times F$, such that the following diagram commutes
[…]
where $\mathrm{pr}_{1}:U\times F\to U$ is the projection onto the first factor. Thus, $(\mathcal{M},\Xi,\mathcal{B})$ is a smooth fiber bundle with fiber $F$.
The state $\Xi(S,\boldsymbol{a},\boldsymbol{\lambda})$ depends only on the $n$ values $\boldsymbol{\mu}(S,\boldsymbol{a},\boldsymbol{\lambda})=(\mu_{1},\dots,\mu_{n})\in\mathbb{R}^{n}$. Hence, $\sigma=\Xi(S,\boldsymbol{a},\boldsymbol{\lambda})$ if and only if $\mu_{i}(S,\boldsymbol{a},\boldsymbol{\lambda})=c_{i}$, where $\boldsymbol{c}=(c_{1},\dots,c_{n})$ determines $\sigma=\rho_{\boldsymbol{c}}$ via the Gibbs form. The fiber $F_{\sigma}$ is therefore the joint level set
$F_{\sigma}=\bigl\{(S,\boldsymbol{a},\boldsymbol{\lambda})\in\mathcal{M}\;\big|\;\mu_{i}(S,\boldsymbol{a},\boldsymbol{\lambda})=c_{i},i=1,\ldots,n\bigr\},$ (24)
a smooth $(n+1)$-dimensional submanifold under the assumptions stated earlier.
Remark 8.
The equilibrium submanifold $\mathcal{E}\subset\mathcal{M}$ is a Legendrian section of the bundle. For $\sigma\in\mathcal{B}$, the intersection $F_{\sigma}\cap\mathcal{E}$ consists of a single element, assuming the map $\boldsymbol{\lambda}\mapsto\rho_{\boldsymbol{\lambda}}$ is injective, corresponding to the unique thermodynamic coordinates $\left(S(\boldsymbol{\lambda}),\boldsymbol{a}(\boldsymbol{\lambda}),\boldsymbol{\lambda}\right)$, where $\mu_{i}=\lambda_{i}$, $S=-\operatorname{\mathrm{tr}}(\rho\ln\rho)$, and $\boldsymbol{a}=\langle\boldsymbol{A}\rangle_{\rho}$. This point satisfies the contact constraint $\eta=0$, i.e.
$dS=\sum_{i=1}^{n}\lambda_{i}da_{i}.$ (25)
Points in $F_{\sigma}\setminus\mathcal{E}$ represent non-equilibrium thermodynamic configurations that still yield the same physical state $\sigma$ but violate equilibrium relations, e.g. inconsistent $\boldsymbol{a}$.
The physical interpretation is that the fiber $F_{\sigma}$ is the set of all thermodynamic labels $(S,\boldsymbol{a},\boldsymbol{\lambda})$ compatible with the same physical density matrix, $\sigma$. On $\mathcal{E}$, we have equilibrium, i.e. unique $S$, correct $\boldsymbol{a}=\langle\boldsymbol{A}\rangle_{\sigma}$, $\boldsymbol{\lambda}$ as intensive parameters. Off $\mathcal{E}$, the points represent non-equilibrium, meaning same $\sigma$, but $S\neq-\operatorname{\mathrm{tr}}(\sigma\ln\sigma)$, or $\boldsymbol{a}\neq\langle\boldsymbol{A}\rangle_{\sigma}$
Example 1.
To build intuition, consider a qubit, with $\mathcal{H}\cong\mathbb{C}^{2}$, so $m=2$, with a single observable $A_{1}=\sigma_{z}$, the Pauli-$Z$ matrix. This is given by
$\sigma_{z}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$. (26)
Then, $\mathcal{M}\cong\mathbb{R}^{3}$ with coordinates $(S,a,\lambda)$; $\mathcal{B}\subset\mathcal{D}^{\circ}$ is the thermal curve of Gibbs states
$\rho_{\lambda}=\frac{1}{2}\begin{pmatrix}1+\tanh\lambda&0\\ 0&1-\tanh\lambda\end{pmatrix}$, (27)
parametrized by inverse temperature $\lambda\in\mathbb{R}$; and the expectation value of the observable and the entropy are
$\begin{split}a(\lambda)&=\langle\sigma_{z}\rangle=\tanh\lambda,\\ S(\lambda)&=\log 2-\frac{1+a(\lambda)}{2}\log(1+a(\lambda))-\frac{1-a(\lambda)}{2}\log(1-a(\lambda)).\end{split}$ (28)
The state function $\Xi:\mathbb{R}^{3}\to\mathcal{B}$ assigns to each thermodynamic point $(S,a,\lambda)$ the density matrix
$\Xi(S,a,\lambda)=\frac{\exp(-\mu(S,a,\lambda)\sigma_{z})}{\operatorname{\mathrm{tr}}[\exp(-\mu(S,a,\lambda)\sigma_{z})]},$ (29)
where $\mu:\mathbb{R}^{3}\to\mathbb{R}$ is any smooth extension of the equilibrium relation $\mu(S(\lambda),a(\lambda),\lambda)=\lambda$.
Geometrically, this means that $\mathcal{M}\cong\mathbb{R}^{3}$ is a 3D volume; $\mathcal{B}\subset\mathcal{D}^{\circ}$ is a 1D curve in the Bloch ball interior; the fiber $F_{\sigma}=\Xi^{-1}(\sigma)$ over a fixed thermal state $\sigma=\rho_{\lambda_{0}}$ is the set of all $(S,a,\lambda)$ such that $\mu(S,a,\lambda)=\lambda_{0}$, i.e. a 2D surface (codimension 1) in $\mathbb{R}^{3}$.
The equilibrium submanifold $\mathcal{E}$ is the curve $\mathcal{E}=\{(S(\lambda),a(\lambda),\lambda)\mid\lambda\in\mathbb{R}\}$, which intersects each fiber $F_{\sigma}$ at exactly one point—the unique thermodynamic coordinates consistent with thermal equilibrium. Thus, multiple points in $\mathcal{M}$ map to the same physical state $\sigma$, but only one lies on $\mathcal{E}$. Points off $\mathcal{E}$ in $F_{\sigma}$ represent non-equilibrium thermodynamic descriptions of $\sigma$: same density matrix, but incorrect entropy $S$, mismatched expectation $a$, or inconsistent intensive parameter $\lambda$. This redundancy is the hallmark of the fiber bundle structure.
Remark 9 (Physical implications of the injectivity and the zeroth law of quantum thermodynamics).
The zeroth law—the transitivity of thermal equilibrium—rests on the injectivity of the Gibbs map $\boldsymbol{\lambda}\mapsto\rho_{\boldsymbol{\lambda}}$.
In the quantum thermodynamic fiber bundle, this injectivity ensures that each Gibbs state $\sigma\in\mathcal{B}$ intersects the equilibrium submanifold $\mathcal{E}$ at exactly one point. Thermal equilibrium between $A$ and $B$ therefore corresponds to both systems mapping to the same fiber $F_{\rho_{A}=\rho_{B}}$, and thus to the same unique equilibrium point $p_{\sigma}\in\mathcal{E}$. The shared coordinates $\boldsymbol{\lambda}$ define a universal intensive parameter field on $\mathcal{E}$, with level sets, e.g. fixed inverse temperature $\beta$, forming transitive equilibrium classes. The zeroth law is thereby elevated from empirical observation to geometric necessity: thermal equilibrium is transitive because equilibrium is unique.
If injectivity fails, by e.g. redundant observables $A_{2}=cA_{1}$, with $c\in\mathbb{R}$, multiple $\boldsymbol{\lambda}$ yield the same $\rho$, destroying uniqueness and allowing inconsistent temperature assignments, the zeroth law breaks. Thus, injectivity is not an additional assumption but the geometric expression of thermodynamic uniqueness: a single quantum state admits one and only one equilibrium thermodynamic description. This structural fact is what allows temperature (and all intensive parameters) to be unambiguously defined, shared across systems in contact, and transitive under thermal equilibration—the very essence of the quantum zeroth law.
VI Movement along fibers
The fibers $F_{\sigma}=\Xi^{-1}(\sigma)$ of the quantum thermodynamic fiber bundle provide a natural arena for analyzing thermodynamic processes at a fixed quantum state $\sigma\in\mathcal{B}$. A path $\gamma:[0,T]\to\mathcal{M}$ confined to $F_{\sigma}$ evolves the thermodynamic coordinates $(S,\boldsymbol{a},\boldsymbol{\lambda})$ while preserving the density operator $\Xi[\gamma(t)]=\sigma$, for all $t\in[0,T]$.
Definition 12.
A smooth path $\gamma:[0,T]\to\mathcal{M}$ is state-preserving with respect to $\sigma\in\mathcal{B}$ if $\gamma(t)\in F_{\sigma}$, for all $t\in[0,T]$, i.e. $\mu_{i}\big(S(t),\boldsymbol{a}(t),\boldsymbol{\lambda}(t)\bigr)=c_{i}$, for $i=1,\dots,n$, where $\boldsymbol{c}\in\mathbb{R}^{n}$ satisfies $\sigma=\rho_{\boldsymbol{c}}$ via the Gibbs form.
The fiber $F_{\sigma}$ is a smooth $(n+1)$-dimensional submanifold of $\mathcal{M}$, defined by $n$ independent level set constraints. Assuming injectivity of $\boldsymbol{\lambda}\mapsto\rho_{\boldsymbol{\lambda}}$, the equilibrium point in $F_{\sigma}$ is the unique intersection $p_{\sigma}:=F_{\sigma}\cap\mathcal{E}$, where $\mu_{i}=\lambda_{i}$, $a_{i}=\operatorname{\mathrm{tr}}(A_{i}\sigma)$, and $S=-\operatorname{\mathrm{tr}}(\sigma\ln\sigma)$. A relaxation process to the equilibrium is a path $\gamma:[0,\infty)\to F_{\sigma}$ with the initial condition $\gamma(0)\in F_{\sigma}\setminus\mathcal{E}$, and limit $\lim_{t\to\infty}\gamma(t)=p_{\sigma}$.
The contact distribution $\ker\eta\subset T\mathcal{M}$ defines Legendre submanifolds of dimension $n$. While $F_{\sigma}$ is $(n+1)$-dimensional, its intersection with $\ker\eta$ gives reversible directions.
Remark 10.
Paths $\gamma(t)\in F_{\sigma}$ keep the density operator $\sigma$ fixed, so the expectation values $a_{i}(t)=\operatorname{\mathrm{tr}}(A_{i}\sigma)$ are target coordinates that may differ from the actual $\operatorname{\mathrm{tr}}(A_{i}\sigma)$ except on the equilibrium point $p_{\sigma}=F_{\sigma}\cap\mathcal{E}$. Thermodynamically, these paths in $F_{\sigma}$ enable computation of quantities like work or heat. The quantum first law can be expressed through changes in the expectation values, constrained by the structure of the fiber. For a quasistatic process, the work done is related to changes in $\boldsymbol{\lambda}(t)$, while heat is associated with $S^{\prime}(t)$. The contact distribution $\ker\eta\subset T\mathcal{M}$ selects the reversible directions inside $F_{\sigma}$. Any deviation from $\ker\eta$ generates positive entropy production, providing a geometric criterion for thermodynamic reversibility. Note that if $\gamma(t)\in\ker\eta$, the process is reversible since there is no entropy production associated to such process.
VII Distances
The preceding discussion on relaxation processes within the fiber suggests that this geometric formulation can quantify the separation between quantum states, whether in equilibrium or non-equilibrium, to measure their distance from the equilibrium configuration on the Legendrian submanifold $\mathcal{E}$.
In quantum theory, the Bures-Wasserstein distance defines a distance measure comparing quantum states, represented by density matrices, by considering the properties of positive-definite self-adjoint matrices [oostrum22_bures]. This distance is particularly suited for the quantum thermodynamic state space, as it aligns with the Riemannian geometry of the space of quantum states $\mathcal{D}$.
Definition 13.
Let $A,B$ be positive-definite self-adjoint matrices on a Hilbert space $\mathcal{H}$. The Bures-Wasserstein distance is defined as
$d_{\mathrm{BW}}(A,B)=\left\{\operatorname{\mathrm{tr}}(A)+\operatorname{\mathrm{tr}}(B)-2\operatorname{\mathrm{tr}}\left[\left(A^{1/2}BA^{1/2}\right)^{1/2}\right]\right\}^{1/2}.$ (30)
Note that, if $\operatorname{\mathrm{tr}}(A)=\operatorname{\mathrm{tr}}(B)=1$, $d_{\mathrm{BW}}(A,B)$ simplifies to
$d_{\mathrm{BW}}(A,B)=\left\{2-2\operatorname{\mathrm{tr}}\left[\left(A^{1/2}BA^{1/2}\right)^{1/2}\right]\right\}^{1/2},$ (31)
where we can define $F(A,B):=\operatorname{\mathrm{tr}}\left[\sqrt{A^{1/2}BA^{1/2}}\right]$ the fidelity between matrices $A$ and $B$. The Bures-Wasserstein distance is a Riemannian distance, inducing a Riemannian metric on the space of quantum states.
Definition 14.
For a tangent vector $X\in T_{\rho_{\boldsymbol{\lambda}_{0}}}\mathcal{B}$ at a Gibbs state $\rho_{\boldsymbol{\lambda}_{0}}\in\mathcal{B}$, the directional derivative of the state map is given by
$X(\rho_{\boldsymbol{\lambda}})=\left.\frac{d}{dt}\right|_{t=0}(\rho_{\boldsymbol{\lambda}}\circ\gamma)(t),$ (32)
where $\gamma:(-\varepsilon,\varepsilon)\longrightarrow\mathcal{B}$ is a smooth curve such that $\gamma(0)=\rho_{\boldsymbol{\lambda}_{0}}$ and $\gamma^{\prime}(0)=X$.
Remark 11.
The above definition can be shown to be independent of the specific curve chosen [caja12_variedades].
Proposition 5.
Let $\rho_{\boldsymbol{\lambda}}\in\mathcal{B}$ be a Gibbs state, and let $X,Y\in T_{\rho_{\boldsymbol{\lambda}}}\mathcal{B}$ be tangent vectors. The Bures-Wasserstein metric is given by
$g_{\mathrm{BW}}(X,Y)=\mathrm{Re}\left[\operatorname{\mathrm{tr}}\left(L_{X}\rho_{\boldsymbol{\lambda}}L_{Y}\right)\right],$ (33)
where $L_{X}$ is the symmetric logarithmic derivative satisfying $\rho_{\boldsymbol{\lambda}}L_{X}+L_{X}\rho_{\boldsymbol{\lambda}}=2X(\rho_{\boldsymbol{\lambda}})$, and similarly for $L_{Y}$. This defines a Riemannian metric on $\mathcal{B}$.
Proof.
The proof is detailed in [oostrum22_bures]. In brief, the metric $g_{\mathrm{BW}}$ arises from the infinitesimal form of the Bures-Wasserstein distance, where $L_{X}$ solves the Lyapunov equation for the perturbation $X(\rho_{\boldsymbol{\lambda}})$. The real part ensures symmetry, and the trace preserves positive-definiteness, establishing $g_{\mathrm{BW}}$ as a Riemannian metric. ∎
In local coordinates, the metric components are
$g_{ij}(\boldsymbol{\lambda})=\textrm{Re}\left[\operatorname{\mathrm{tr}}\left(\rho_{\boldsymbol{\lambda}}L_{i}L_{j}\right)\right],\quad L_{i}=\frac{\partial\ln(\rho_{\boldsymbol{\lambda}})}{\partial\lambda_{i}},$ (34)
with the line element
$ds^{2}=\sum_{i,j=1}^{n}g_{ij}d\lambda^{i}d\lambda^{j}.$ (35)
Corollary 1.
$(\mathcal{B},g_{\mathrm{BW}})$ is a Riemannian manifold.
The Riemannian structure of $(\mathcal{B},g_{\mathrm{BW}})$ is further supported by the following fundamental result.
Theorem 3 (Hopf-Rinow).
For a connected Riemannian manifold $(M,g)$, the following are equivalent:
  1. 1.
    $M$ is complete as a metric space.
  2. 2.
    $M$ is geodesically complete.
  3. 3.
    A subset in $M$ is compact if and only if it is closed and bounded.
Additionally, if any of the above holds, then any two points $p,q\in M$ can be joined by a minimizing geodesic.
VIII Geodesics and quasistatic processes
The Bures-Wasserstein metric $g_{\mathrm{BW}}$ endows the submanifold of Gibbs states $\mathcal{B}$ with a Riemannian structure, enabling the geometric framework for analyzing quasistatic thermodynamic processes.
Definition 15.
Let $(\mathcal{B},g_{\mathrm{BW}})$ be the Riemannian manifold of Gibbs states. Let $\boldsymbol{\lambda}(t)=(\lambda_{1}(t),\dots,\\ \lambda_{n}(t))$ be a curve in coordinate space. The corresponding path in $\mathcal{B}$ is then $\gamma(t)=\rho_{\boldsymbol{\lambda}(t)}:[0,T]\to\mathcal{B}$. The thermodynamic length $L(\gamma)$ is
$L(\gamma)=\int_{0}^{T}\sqrt{g_{\mathrm{BW}}(\gamma^{\prime}(t),\gamma^{\prime}(t))}dt=\int_{0}^{T}\sqrt{\sum_{i,j=1}^{n}g_{ij}(\boldsymbol{\lambda}(t))\lambda_{i}^{\prime}(t)\lambda_{j}^{\prime}(t)}dt,$ (36)
where $\gamma^{\prime}(t)=\sum_{i=1}^{n}\lambda_{i}^{\prime}(t)\partial_{\lambda^{i}}\big|_{\gamma(t)}\in T_{\gamma(t)}\mathcal{B}$ is the pushforward of the coordinate velocity, and $\lambda_{i}^{\prime}(t)=d\lambda_{i}(t)/dt$, for $i=1,\ldots,n$.
The Bures-Wasserstein distance quantifies the distinguishability between Gibbs states $\rho_{\boldsymbol{\lambda}}$, encoding quantum fluctuations in response to changes in control parameters. Paths $\gamma(t)$ in $\mathcal{B}$ correspond to sequences of instantaneous equilibrium states.
A quasistatic process varies $\lambda(t)$ sufficiently slowly so that the system remains in the Gibbs state $\rho_{\boldsymbol{\lambda}}$ at each instant. Such processes are reversible, producing zero entropy production. Minimizing geodesics on $(\mathcal{B},g_{\mathrm{BW}})$, which assuming completeness of the Riemannian manifold $(\mathcal{B},g_{\mathrm{BW}})$ exist by virtue of the Hopf-Rinow theorem, thus represent optimal quasistatic transformations, evolving the system through Gibbs states while minimizing the thermodynamic length $L(\gamma)$. This length quantifies cumulative state change and bounds the minimal work required to drive the transformation.
Definition 16.
The entropy production rate along a finite-speed path $\gamma:[0,T]\to\mathcal{B}$ is
$\varsigma_{\gamma}(t)=\kappa g_{\mathrm{BW}}(\gamma^{\prime}(t),\gamma^{\prime}(t))=\kappa\sum_{i,j=1}^{n}g_{ij}(\boldsymbol{\lambda}(t))\lambda_{i}^{\prime}(t)\lambda_{j}^{\prime}(t),$ (37)
where $\kappa$ is a system-dependent constant ensuring $\varsigma(t)$ has units of entropy per time. Total entropy production along $\gamma(t)$ is
$\Sigma_{\gamma}(t)=\int_{0}^{T}\varsigma(t)dt=\int_{0}^{T}\kappa g_{\mathrm{BW}}(\gamma^{\prime}(t),\gamma^{\prime}(t))dt.$ (38)
For finite-speed processes, $\varsigma_{\gamma}(t)>0$ due to non-zero velocity $\gamma^{\prime}(t)$, indicating irreversibility [tejero2025entropy]. In the quasistatic limit ($\lambda_{i}^{\prime}\to 0$), $\varsigma_{\gamma}(t)\to 0$ and $\Sigma_{\gamma}(t)\to 0$. Minimizing geodesics minimize both $L(\gamma)$ and $\Sigma$, optimizing the path to reduce the entropy production, thus achieving maximal reversibility.
Remark 12.
The term $g_{\mathrm{BW}}(\gamma^{\prime},\gamma^{\prime})$ measures the instantaneous rate of state evolution, driving irreversibility via excitations or non-equilibrium effects. Minimizing geodesics mitigate these by following the smoothest path in state space, allowing the system to remain in equilibrium at each step, analogous to adiabatic transformations. The proportionality $\varsigma_{\gamma}\propto g_{\mathrm{BW}}(\gamma^{\prime},\gamma^{\prime})$ reflects that faster transformations disrupt equilibrium more severely, increasing $\Sigma_{\gamma}$.
The following results formalize this connection between geodesics and quasistatic processes.
Proposition 6.
Let $(\mathcal{B},g_{\mathrm{BW}})$ be the Riemannian manifold of equilibrium states. Optimal quasistatic quantum thermodynamic processes are minimizing geodesics in $(\mathcal{B},g_{\mathrm{BW}})$.
Proof.
Quasistatic processes evolve through Gibbs states $\rho_{\boldsymbol{\lambda}}$ with minimal dissipation. The Bures-Wasserstein metric quantifies state distinguishability, and minimizing geodesics minimize thermodynamic length $L(\gamma)$, corresponding to paths of least entropy production $\Sigma_{\gamma}$, as required for optimal quasistatic transformations. ∎
Theorem 4 (Geodesic connectivity).
Let $(\mathcal{B},g_{\mathrm{BW}})$ be a complete Riemannian manifold. Then any two equilibrium states $\rho_{\boldsymbol{\lambda}_{1}},\rho_{\boldsymbol{\lambda}_{2}}\in\mathcal{B}$ are connected by a minimizing geodesic, representing an optimal quasistatic transformation.
Proof.
This result is a direct consequence of the Hopf-Rinow theorem. Since $(\mathcal{B},g_{\mathrm{BW}})$ is a Riemannian manifold, by the Hopf-Rinow theorem completeness of $(\mathcal{B},g_{\mathrm{BW}})$ implies that any two points can be joined by a minimizing geodesic. ∎
The Hopf-Rinow theorem ensures that $\mathcal{B}$ is a globally accessible manifold for thermodynamic transformations, with minimizing geodesics providing the optimal paths for quasistatic processes, minimizing both thermodynamic length and entropy production.
There might be situations where the curve $\gamma$ in $\mathcal{B}$ does not necessarily represent a geodesic curve. Such curves, deviating from these shortest paths, describe quasistatic processes with redundant or oscillatory changes in $\boldsymbol{\lambda}$, such as an isothermal process with inefficient parameter adjustments, yet still preserving the Gibbs state. These non-optimal paths incur higher thermodynamic costs, reflecting practical constraints or suboptimal control in quantum protocols, like those in thermal machines or quantum control. For example, an isothermal transformation maintaining $\rho_{\beta}(t)$ at all times is quasistatic (always in equilibrium), but only the geodesics can minimize dissipation.
IX Boundary effects
Proposition 7.
Let $\mathcal{D}^{\circ}$ be the $(m^{2}-1)$-dimensional smooth manifold of full-rank density operators on a finite-dimensional Hilbert space $\mathcal{H}$ with $\dim\mathcal{H}=m$, embedded as the interior of the compact convex set $\mathcal{D}$ of all density operators. Then:
  1. 1.
    For a state $\rho\in\mathcal{D}^{\circ}$, with eigenvalues $p_{i}$ with $i=1,\dots,m$ (not necessarily all distinct), the entropy extends continuously from $\mathcal{D}^{\circ}$ to the boundary $\partial\mathcal{D}$.
  2. 2.
    For any smooth curve $\gamma:[0,1)\to\mathcal{D}^{\circ}$ such that $\lim_{t\to 1^{-}}\gamma(t)=\overline{\rho}\in\partial\mathcal{D}$ with $\mathrm{rank}(\overline{\rho})=k<m$, the entropy satisfies
    $\lim_{t\to 1^{-}}S(\gamma(t))=S(\overline{\rho}),$ (39)
    where $S(\overline{\rho})=-\sum_{i=1}^{k}\overline{p}_{i}\ln\overline{p}_{i}$, and $\overline{p}_{i}$ with $i=1,\dots,k$ (not necessarily all distinct) are the positive eigenvalues of $\overline{\rho}$, where $k$ is the rank of $\overline{\rho}$. In particular, if $\overline{\rho}$ is a pure state $(k=1)$, then $S(\gamma(t))\to 0$, and if $\overline{\rho}$ is maximally mixed on its support, i.e. $\overline{p}_{i}=1/k$ for $i=1,\ldots,k$, then $S(\gamma(t))\to\ln k$.
Proof.
First, we establish the continuous extension of $S$ to $\partial\mathcal{D}$. For a state $\overline{\rho}\in\mathcal{D}$ with eigenvalues $p_{i}\geq 0$, some possibly zero, the entropy is defined by extending the function $f(x)=-x\ln x$ to the domain boundary $x=0$. Let $\gamma:[0,1)\to\mathcal{D}^{\circ}$ be a smooth curve with $\lim_{t\to 1^{-}}\gamma(t)=\overline{\rho}\in\partial\mathcal{D}$ of rank $k<m$. In a neighborhood of $\overline{\rho}$, $\gamma(t)$ can be diagonalized: $\gamma(t)=U(t)\operatorname{diag}(p_{1}(t),\ldots,p_{m}(t))U^{*}(t)$ with $p_{i}(t)>0$, $\sum_{i=1}^{m}p_{i}(t)=1$, and $U(t)$ is a $m\times m$ unitary matrix for each value of $t$. As $t\to 1^{-}$, assume without loss of generality that $p_{i}(t)\to\overline{p}_{i}>0$ for $i=1,\ldots,k$, and $p_{i}(t)\to 0$ for $i=k+1,\ldots,m$, with $\sum_{i=1}^{k}\overline{p}_{i}=1$. Then
$S(\gamma(t))=-\sum_{i=1}^{k}p_{i}(t)\ln p_{i}(t)-\sum_{i=k+1}^{m}p_{i}(t)\ln p_{i}(t).$ (40)
The first sum converges to $S(\overline{\rho})$ by continuity of $f(x)$ at $\overline{p}_{i}>0$. For the second sum, consider $f(p_{i}(t))=-p_{i}(t)\ln p_{i}(t)$. As $p_{i}(t)\to 0^{+}$, $f(p_{i}(t))\to 0$. The sum $\sum_{i=k+1}^{m}f(p_{i}(t))$ has mkm-k terms and each term converges to 0, so $\sum_{i=k+1}^{m}p_{i}(t)\ln p_{i}(t)\to 0$. Thus,
$\lim_{t\to 1^{-}}S(\gamma(t))=S(\overline{\rho}).$ (41)
The special cases follow immediately: pure states give $S(\overline{\rho})=0$; maximal mixing on support gives $S(\overline{\rho})=\ln k$. ∎
Note that $\partial\mathcal{D}$ is stratified by rank, where the rank-$k$ stratum is a smooth manifold. The stratification ensures that $\partial\mathcal{D}$ is decomposed into smooth manifolds (strata), each corresponding to a fixed rank $k$. The entropy $S(\rho)$ for $\rho$ in the rank-$k$ stratum depends on the eigenvalue distribution.
Theorem 5 (Geometric unattainability of the boundary).
Let $(\mathcal{B},g_{\mathrm{BW}})$ be the Riemannian manifold of Gibbs states. Then no geodesic $\gamma:[0,T]\to\mathcal{B}$ of finite length can reach $\partial\mathcal{D}$.
Proof.
Consider a boundary element described by a state $\rho\in\partial\mathcal{D}$ with $\mathrm{rank}(\rho)<m$. A geodesic curve $\gamma(t):[0,1)\longrightarrow\mathcal{B}$ such that $\lim_{t\to 1^{-}}\gamma(t)=\rho$ has eigenvalues $p_{i}(t)$ with at least one $p_{k}(t)\to 0^{+}$ under such limit. Using the Bures-Wasserstein metric, the line element $ds^{2}$ is seen to contain a term proportional to $p_{k}^{-1}$ when evaluated over $\gamma$. Since the resulting integral is divergent, $L(\gamma)$ diverges. ∎
A thermodynamic process in the quantum thermodynamic state space $\mathcal{M}$ projects to a curve $\gamma$ in $\mathcal{D}^{\circ}$ via $\Xi$. Reaching boundary states in $\partial\mathcal{D}$ would require infinite thermodynamic length, hence infinite resources.
Corollary 2 (Third law of quantum thermodynamics).
Let $\gamma:[0,1)\longrightarrow\mathcal{B}$ be a smooth curve such that $\lim_{t\longrightarrow 1^{-}}\gamma(t)\in\partial\mathcal{D}$ is a maximally mixed state. Then, the entropy function $S:\mathcal{D}^{\circ}\to\mathbb{R}^{+}$ satisfies
$\lim_{t\longrightarrow 1^{-}}S(\gamma(t))=\ln k,$ (42)
where $k$ is the rank of the target boundary stratum. Moreover, boundary strata, are unattainable along any finite-length geodesic in $\mathcal{B}$.
Proof.
From the theorem, no finite-length geodesic can reach $\partial\mathcal{D}$. By the proposition, $S(\gamma(t))\to\ln k$, as $S$ is continuous on $\mathcal{B}$. For pure states, $k=1$, meaning $S\to 0$ in infinite time. ∎
Corollary 3.
There exists no finite-length thermodynamic process transforming a full-rank Gibbs state into a pure state.
Remark 13.
The classical third law states that absolute zero temperature is unattainable in finite steps. Here, zero-entropy states (pure states) play an analogous role, but the unattainability arises purely from the Riemannian geometry of the state space, not from any explicit parameter. This geometric third law holds for any control protocol in $\mathcal{M}$ projecting to $\mathcal{D}^{\circ}$.
X Connections, curvature and holonomy in the quantum thermodynamic fiber bundle
The fiber $F_{\sigma}=\Xi^{-1}(\sigma)$ contains all thermodynamic configurations yielding the fixed quantum state $\sigma\in\mathcal{D}$, with only the equilibrium point $p_{\sigma}=F_{\sigma}\cap\mathcal{E}$ lying on the Legendrian submanifold $\mathcal{E}$. Geodesics on $\mathcal{E}$ connect equilibrium states across different fibers, while non-equilibrium points in $F_{\sigma}\setminus\mathcal{E}$ represent transient configurations. Relaxation paths within $F_{\sigma}$ converge to $p_{\sigma}$, and their dissipation can be quantified and minimized using a pseudo-Riemannian metric $g_{\mathcal{M}}$ on $\mathcal{M}$ that extends the Bures-Wasserstein metric $g_{\mathrm{BW}}$ from $\mathcal{E}$.
Thermodynamic processes, either quasistatic evolutions along Gibbs states or driven non-equilibrium dynamics, correspond to paths in $\mathcal{M}$ projecting via $\Xi$ to curves in $\mathcal{D}$. To model these transitions geometrically, we introduce an Ehresmann connection on the fiber bundle, decomposing $T\mathcal{M}$ into vertical and horizontal subbundles. This connection enables parallel transport of thermodynamic states along paths in $\mathcal{D}$, with curvature quantifying non-integrability and holonomy inducing geometric irreversibility in cyclic processes. This is analogous to holonomies existent in gauge theories.
Definition 17.
The vertical subbundle $\mathscr{V}\subset T\mathcal{M}$ is
$\mathscr{V}_{p}=\ker\left(d\Xi_{p}:T_{p}\mathcal{M}\to T_{\Xi(p)}\mathcal{D}\right)=T_{p}F_{\Xi(p)},$ (43)
the tangent space to the fiber $F_{\sigma}$ at $p\in F_{\sigma}$. Since $\dim\mathcal{M}=2n+1$ and $\dim\mathcal{D}=n$, we have $\dim\mathscr{V}_{p}=n+1$. Assuming that $\Xi$ is a submersion, $d\Xi_{p}$ is surjective, and $\mathscr{V}=\ker(d\Xi)$ is a smooth subbundle.
Definition 18.
An Ehresmann connection is a smooth horizontal subbundle $\mathscr{H}\subset T\mathcal{M}$ such that
  1. 1.
    $T_{p}\mathcal{M}=\mathscr{H}_{p}\oplus\mathscr{V}_{p}$ for all $p\in\mathcal{M}$,
  2. 2.
    $d\Xi_{p}|_{\mathscr{H}_{p}}:\mathscr{H}_{p}\to T_{\Xi(p)}\mathcal{D}$ is a linear isomorphism.
The dimension of this subbundle is $\dim\mathscr{H}_{p}=n$.
Let $\pi:\mathcal{M}\to\mathcal{E}$ be the projection along fibers to the equilibrium submanifold. At any point $p=(S,\boldsymbol{a},\boldsymbol{\lambda})\in\mathcal{M}$, the tangent space is spanned by
$T_{p}\mathcal{M}=\operatorname{span}\left\{\frac{\partial}{\partial S},\ \frac{\partial}{\partial a_{1}},\ \dots,\ \frac{\partial}{\partial a_{n}},\ \frac{\partial}{\partial\lambda_{1}},\ \dots,\ \frac{\partial}{\partial\lambda_{n}}\right\}.$ (44)
For $p\in F_{\sigma}$ with $\mu_{i}=\lambda_{i}$, we have $\lambda_{i}=c_{i}$, for $i=1,\ldots,n$, so
$\mathscr{V}_{p}=T_{p}F_{\sigma}=\operatorname{span}\left\{\frac{\partial}{\partial S},\ \frac{\partial}{\partial a_{1}},\ \dots,\ \frac{\partial}{\partial a_{n}}\right\},$ (45)
and the subbundle
$\mathscr{H}_{p}=\operatorname{span}\left\{\frac{\partial}{\partial\lambda_{1}},\ \dots,\ \frac{\partial}{\partial\lambda_{n}}\right\}$ (46)
projects isomorphically to $T_{p_{\sigma}}\mathcal{E}$ via $\pi_{*}$.
Proposition 8.
Consider a pseudo-Riemannian metric $g_{\mathcal{M}}$ on $\mathcal{M}$ such that $g_{\mathcal{M}}|_{\mathscr{V}_{p}}$ is positive-definite. This metric induces an Ehresmann connection by defining
$\mathscr{H}_{p}=\{v\in T_{p}\mathcal{M}\mid g_{\mathcal{M}}(v,w)=0,\ \forall w\in\mathscr{V}_{p}\}.$ (47)
Proof.
For any tangent vector $u\in T_{p}\mathcal{M}$, decompose $u=h+v$, with $h\in\mathscr{H}_{p}$, $v\in\mathscr{V}_{p}$. The condition $g_{\mathcal{M}}(h,w)=0$ for all $w\in\mathscr{V}_{p}$ determines $v$ via the linear system
$g_{\mathcal{M}}(u-v,w_{\beta})=0,$ (48)
where $\beta=1,\ldots,n+1$, for a basis $\{w_{\beta}\}$ of $\mathscr{V}_{p}$. The matrix $g_{\mathcal{M}}(w_{\alpha},w_{\beta})$ is invertible since $g_{\mathcal{M}}|_{\mathscr{V}_{p}}$ is positive-definite, ensuring a unique $v$.
For the second condition in Definition 18, since $\mathscr{V}_{p}=\ker d\Xi_{p}$,
$d\Xi_{p}(T_{p}\mathcal{M})=d\Xi_{p}(\mathscr{H}_{p}).$ (49)
Given $\dim\mathscr{H}_{p}=\dim T_{\Xi(p)}\mathcal{D}=n$ and $\Xi$ a submersion, $d\Xi_{p}|_{\mathscr{H}_{p}}$ is an isomorphism. The smoothness of $\mathscr{H}$ follows from the smoothness of $g_{\mathcal{M}}$ and $\Xi$. ∎
Remark 14.
In order to define such a metric $g_{\mathcal{M}}$, the desiderata are:
  1. 1.
    $g_{\mathcal{M}}|_{T\mathcal{E}}=g_{\mathrm{BW}}$,
  2. 2.
    compatibility with the contact structure $\eta$,
  3. 3.
    positive-definiteness on $\ker(\eta)$ and controlled signature in transverse directions.
Proposition 9.
A pseudo-Riemannian metric on $\mathcal{M}$, $g_{\mathcal{M}}$, satisfying the desiderata in Remark 14 is
$g_{\mathcal{M}}=g_{S}dS^{2}+\sum_{i=1}^{n}g_{a_{i}}da_{i}^{2}+\pi^{*}g_{\mathrm{BW}}+\sum_{i=1}^{n}h_{i}(dS\otimes d\lambda_{i}+d\lambda_{i}\otimes dS),$ (50)
where:
  • $g_{S}(p_{\sigma})\in\mathbb{R}$ controls the entropy direction, which may be negative,
  • $g_{a_{i}}(p_{\sigma})>0$ are positive-definite on expectation value directions, for all $i=1,\ldots,n$,
  • $\pi^{*}g_{\mathrm{BW}}=\sum_{i,j}g_{ij}(\pi(p_{\sigma}))d\lambda_{i}d\lambda_{j}$ pulls back the Bures-Wasserstein metric,
  • $h_{i}(p)$ are cross-terms ensuring contact compatibility, for all $i=1,\ldots,n$.
Note that, on $\mathcal{E}$, $\eta=0$ implies $dS=\sum\lambda_{i}da_{i}$, so
$g_{\mathcal{M}}|_{T\mathcal{E}}=\pi^{*}g_{\mathrm{BW}}.$ (51)
On a fiber $F_{\sigma}$, where $\lambda_{i}=c_{i}$ fixed, $d\lambda_{i}=0$ for all $i=1,\ldots,n$, and
$g_{\mathcal{M}}|_{F_{\sigma}}=g_{S}dS^{2}+\sum_{i=1}^{n}g_{a_{i}}da_{i}^{2}.$ (52)
For a path $\gamma:[0,T]\to\mathcal{M}$ with $\gamma(t)=(S(t),\boldsymbol{a}(t),\boldsymbol{\lambda}(t))$, the thermodynamic length is
$L(\gamma)=\int_{0}^{T}\sqrt{g_{\mathcal{M}}(\gamma^{\prime}(t),\gamma^{\prime}(t))}dt,$ (53)
where
$g_{\mathcal{M}}(\gamma^{\prime},\gamma^{\prime})=g_{S}S^{\prime 2}+\sum_{i=1}^{n}g_{a_{i}}(a_{i}^{\prime})^{2}+\sum_{i,j=1}^{n}g_{ij}\left(\pi(\gamma(t))\right)\lambda_{i}^{\prime}\lambda_{j}^{\prime}+2\sum_{i=1}^{n}h_{i}S^{\prime}\lambda_{i}^{\prime}.$ (54)
Along a relaxation path $\gamma(t)\in F_{\sigma}$, with $\lambda_{i}^{\prime}=0$, for all $i=1,\ldots,n$,
$L(\gamma)=\int_{0}^{T}\sqrt{g_{S}S^{\prime 2}+\sum_{i=1}^{n}g_{a_{i}}(a_{i}^{\prime})^{2}}dt.$ (55)
This represents the thermodynamic length for a non-equilibrium point $\gamma(0)$ in the fiber out of the equilibrium submanifold, toward $\gamma(T)\in F_{\sigma}\cap\mathcal{E}$. In such process, the entropy production rate is
$\varsigma(t)=\kappa g_{\mathcal{M}}(\gamma^{\prime}(t),\gamma^{\prime}(t)),$ (56)
with total
$\Sigma=\kappa\int_{0}^{T}g_{\mathcal{M}}(\gamma^{\prime}(t),\gamma^{\prime}(t))dt,$ (57)
for $\kappa>0$.
To ensure the compatibility with the contact structure, we need the metric $g_{\mathcal{M}}$ to be positive-definite on $\ker(\eta)$, and possibly indefinite in the transverse direction (e.g., along $\partial S$). The Reeb vector field $\mathcal{R}$ is $\mathcal{R}=\partial_{S}$, since $\eta(\mathcal{R})=1$ and $d\eta=-\sum_{i=1}^{n}d\lambda_{i}\wedge da_{i}$. The metric can assign $g_{\mathcal{M}}(\mathcal{R},\mathcal{R})=g_{S}$, with cross-terms $h_{i}$ ensuring that $g_{\mathcal{M}}$ respects the structure of the contact distribution via $d\eta$.
The Ehresmann connection enables parallel transport of thermodynamic states.
Definition 19.
For a smooth path $\gamma:[0,1]\to\mathcal{D}$, with $\gamma(0)=\sigma_{1}$, $\gamma(1)=\sigma_{2}$, and $p_{0}\in\Xi^{-1}(\sigma_{1})$, the horizontal lift is a curve $\tilde{\gamma}:[0,1]\to\mathcal{M}$ satisfying
$\Xi\circ\tilde{\gamma}(t)=\gamma(t),\quad\tilde{\gamma}(0)=p_{0},\quad\tilde{\gamma}^{\prime}(t)\in\mathscr{H}_{\tilde{\gamma}(t)}.$ (58)
The lift is unique due to the second condition in Definition 18. Locally, the lift is determined by the horizontal vector fields of the connection. From the connection, it is straightforward to define the curvature form, which measures the non-integrability of $\mathscr{H}$.
Definition 20.
For horizontal vector fields $X,Y\in\mathfrak{X}(\mathscr{H})$, the curvature is
$R(X,Y)=\operatorname{\mathrm{proj}}_{\mathscr{V}}([X,Y])\in\mathfrak{X}(\mathscr{V}),$ (59)
where $[X,Y]$ is the Lie bracket and $\operatorname{\mathrm{proj}}_{\mathscr{V}}:T\mathcal{M}\to\mathscr{V}$ is the projection onto the vertical subbundle.
The curvature is a $\mathscr{V}$-valued two-form, expressed via the covariant derivative of the connection $\nabla$: $R(X,Y)=\nabla_{X}Y-\nabla_{Y}X-[X,Y]$. In a local trivialization of the bundle, let $\mathcal{D}$ have coordinates $\{b_{k}\}_{k=1}^{n}$, and fibers have coordinates $\{v_{\alpha}\}_{\alpha=0}^{n}$, so $\mathcal{M}$ has coordinates $(b_{k},v_{\alpha})$, and $\Xi(b_{k},v_{\alpha})=\sigma(b_{k})$ The vertical subbundle is $\mathscr{V}=\mathrm{span}\{\partial_{v_{\alpha}}\}_{\alpha=0}^{n}$. Horizontal vector fields are
$e_{k}=\dfrac{\partial}{\partial b_{k}}-\sum_{\alpha=0}^{n}\Gamma_{b_{k}}^{v_{\alpha}}\dfrac{\partial}{\partial v_{\alpha}},$ (60)
where $\Gamma_{b_{k}}^{v_{\alpha}}$ are connection coefficients satisfying $g_{\mathcal{M}}(e_{k},\partial_{v_{\beta}})=0$, for all $\beta=0,\ldots,n$. The curvature components are
$R(e_{k},e_{l})=\operatorname{\mathrm{proj}}_{\mathscr{V}}([e_{k},e_{l}]).$ (61)
The Lie bracket gives the general expression
$\begin{split}[e_{k},e_{l}]&=\left[\dfrac{\partial}{\partial b_{k}}-\sum_{\alpha}\Gamma_{b_{k}}^{v_{\alpha}}\dfrac{\partial}{\partial v_{\alpha}},\dfrac{\partial}{\partial b_{l}}-\sum_{\alpha}\Gamma_{b_{l}}^{v_{\alpha}}\dfrac{\partial}{\partial v_{\alpha}}\right]\\ &=\sum_{\alpha}\left(\dfrac{\partial}{\partial b_{k}}\Gamma_{b_{l}}^{v_{\alpha}}-\dfrac{\partial}{\partial b_{l}}\Gamma_{b_{k}}^{v_{\alpha}}+\sum_{\beta}[\Gamma_{b_{k}}^{v_{\beta}},\Gamma_{b_{l}}^{v_{\alpha}}]\right)\dfrac{\partial}{\partial v_{\alpha}},\end{split}$ (62)
where the commutator is in the Lie algebra of the fiber structure group. The curvature two-form is
$R=\sum_{\alpha}\sum_{k,l}R_{b_{k}b_{l}}^{v_{\alpha}}db_{k}\wedge db_{l}\otimes\dfrac{\partial}{\partial v_{\alpha}}.$ (63)
Holonomy arises when parallel transporting along a closed loop in $\mathcal{D}$.
Definition 21.
For a closed loop $\gamma:S^{1}\to\mathcal{D}$ based at $\sigma\in\mathcal{D}$, with horizontal lift $\tilde{\gamma}:S^{1}\to\mathcal{M}$ starting at $p_{0}\in\Xi^{-1}(\sigma)$, the holonomy is the vertical displacement
$\mathrm{Hol}(\gamma,p_{0})=\tilde{\gamma}(1)-p_{0}\in\mathscr{V}_{p_{0}},$ (64)
where $\tilde{\gamma}(1)\in\Xi^{-1}(\sigma)$.
The idea is that for a closed loop $\gamma$ in $\mathcal{D}$, the horizontal lift starting at $p_{0}\in\Xi^{-1}(\sigma)$ may end at a different point $p_{1}\equiv\tilde{\gamma}(1)\in\Xi^{-1}(\sigma)$, with the displacement $h=p_{1}-p_{0}\in\mathscr{V}_{p_{0}}$. The holonomy is an element of the fiber structure group, acting as a translation in the thermodynamic variables.
Proposition 10.
For a closed loop $\gamma$ bounding a surface $\mathscr{S}\subset\mathcal{D}$, the holonomy is
$\mathrm{Hol}(\gamma,p_{0})=\mathcal{P}\exp\left(-\int_{\mathscr{S}}R\right),$ (65)
where the integral is the fiber-valued integral of the curvature two-form over $\mathscr{S}$.
Proof.
The connection one-form $\omega\in\Lambda^{1}(\mathcal{D})$ satisfies $\nabla_{X}Y=[X,Y]-\omega(X)Y$ for horizontal vector fields $X,Y.$ The curvature is $R=d\omega+\omega\wedge\omega$. For a loop $\gamma$, the holonomy is the path-ordered exponential
$\mathrm{Hol}(\gamma,p_{0})=\mathcal{P}\exp\left(-\oint_{\gamma}\omega\right).$ (66)
where $\mathcal{P}$ denotes path ordering along $\gamma$, and $\omega$ is the $\mathfrak{g}$-valued connection 1-form. For abelian structure groups (e.g. translations in $\mathbb{R}^{n+1}$), this simplifies to
$\mathrm{Hol}(\gamma,p_{0})=\exp\left(-\oint_{\gamma}\omega\right)=\exp\left(-\int_{\mathscr{S}}d\omega\right)=\exp\left(-\int_{\mathscr{S}}R\right),$ (67)
by Stokes theorem, assuming $\omega\wedge\omega=0$. For later interest, in the quantum thermodynamic fiber, the fiber is diffeomorphic to $\mathbb{R}^{n+1}$, with abelian translations, so the result applies. ∎
Remark 15.
The path ordering is essential since $R(t_{1}),R(t_{2})]\neq 0$ in general, for $t_{1}\neq t_{2}$, for all $t_{1},t_{2}\in S^{1}$. The result depends on the homotopy class of $\gamma$ within $\pi_{1}(\mathcal{B})$.
Even in the non-abelian case, holonomy represents a geometric phase acquired during a cyclic evolution of the intensive parameters $\boldsymbol{\lambda}$. Upon returning to the same quantum state $\sigma$, the thermodynamic labels $(S,\boldsymbol{a})$ are transformed by a non-trivial group element $h\neq e$, for all $h\in G$, and $e$ the identity element of $G$. This transformation is not removable by local reparameterization and induces irreversible entropy production when the system is forced to return to its initial thermodynamic state.
We shall see in the following section that the fiber is $\mathbb{R}^{n+1}$ with additive group structure, which is abelian. The non-abelian generalization would arise if the thermodynamic labels were subject to a non-commutative redundancy (e.g., in systems with internal symmetries or constrained ensembles), but such extensions lie beyond the current scope.
Proposition 11.
If the connection coefficients vanish in a trivialization, then $R=0$, and the holonomy of any closed loop is zero.
Proof.
If the connection coefficients $\Gamma=0$, horizontal vector fields are $X=\sum_{k}X^{k}\partial_{b_{k}}$. The Lie bracket is
$X,Y]=\sum_{k,l}\left(X^{k}\dfrac{\partial}{\partial b_{k}}Y^{l}-Y^{k}\dfrac{\partial}{\partial b_{k}}X^{l}\right)\dfrac{\partial}{\partial b_{l}},$ (68)
which lies in $\mathscr{H}$, so $\operatorname{\mathrm{proj}}_{\mathscr{V}}([X,Y])=0$, hence $R=0$. The connection one-form $\omega=0$, so $d\omega=0$, and
$\mathrm{Hol}(\gamma,p_{0})=\exp\left(-\oint_{\gamma}\omega\right)=0.$ (69)
Some corollaries of particular relevance for quantum thermodynamics are the following:
Corollary 4.
Any cyclic thermodynamic process corresponding to a closed loop $\gamma$ in $\mathcal{D}$ with non-zero holonomy $\mathrm{Hol}(\gamma,p_{0})\neq 0$ induces geometric irreversibility, contributing to net entropy production $\Sigma_{\gamma}>0$, unless the connection is flat, reflecting irreversibility due to curvature.
Corollary 5.
If the connection is flat $(R=0)$, cyclic processes in $\mathcal{D}$ are geometrically reversible: the horizontal lift of $\gamma$ returns to the initial point in the fiber, with zero geometric entropy production $\Sigma_{\gamma}=0$.
This means cyclic processes (loops in parameter space like temperature) may return the quantum state but shift thermodynamic variables (e.g. entropy), leading to irreversibility—requiring dissipation to correct the shift.
Remark 16.
The holonomy $\mathrm{Hol}(\gamma,p_{0})$ is a functional not of the specific path $\gamma:S^{1}\to\mathcal{B}$, but of its homotopy class $\gamma]\in\pi_{1}(\mathcal{B},\sigma)$. That is, if $\gamma\sim\gamma^{\prime}$ via a homotopy fixing the base point $\sigma$, then $\mathrm{Hol}(\gamma,p_{0})=\mathrm{Hol}(\gamma^{\prime},p_{0})$. This follows from the horizontal nature of the lift: any deformation of $\gamma$ within $\mathcal{B}$ induces a corresponding deformation of $\tilde{\gamma}$ within $\mathcal{M}$ that remains horizontal, hence returns to the same fiber point after closure.
In the abelian case, this implies that holonomy defines a group homomorphism $\Phi:\pi_{1}(\mathcal{B})\to G$, where $\Phi([\gamma])=h(\gamma,p_{0})$, independent of base point $p_{0}$, since vertical translations commute. The image of $\Phi$ is generated by curvature integrals over a basis of 2-cycles in $H_{2}(\mathcal{B})$, via de Rham cohomology.
In the non-abelian case, holonomy defines a representation of the fundamental group $\rho^{G}:\pi_{1}(\mathcal{B})\to G$, where conjugacy classes of $\rho^{G}([\gamma])$ are base-point independent, and the full holonomy group is the subgroup generated by such elements under concatenation of loops. Non-commutativity implies that the order of traversal matters: $\rho^{G}([\gamma_{1}\cdot\gamma_{2}])\neq\rho^{G}([\gamma_{1}])\rho^{G}([\gamma_{2}])$ in general.
Two cyclic protocols in intensive parameter space $\boldsymbol{\lambda}(t)$ that return to the same Gibbs state $\sigma$ but traverse homotopically distinct paths accumulate different geometric entropy shifts, even in the quasistatic limit. This constitutes a topological contribution to irreversibility—unremovable by slowing down the process—and sets a fundamental lower bound on dissipation in cyclic processes operating over topologically non-trivial control manifolds. The topological irreversibility is then tied to the concept of topological entropy production.
In the current framework, where $\mathcal{B}$ is contractible in many finite-dimensional models (e.g., full-rank Gibbs states), $\pi_{1}(\mathcal{B})=0$, and holonomy vanishes for contractible loops. However, in constrained ensembles, $\pi_{1}(\mathcal{B})$ may be non-trivial, opening the door to topological quantum thermodynamics.
XI Principal Bundle Structure
XI.1 Principal bundle
To elevate the quantum thermodynamic fiber bundle $(\mathcal{M},\mathcal{B},\Xi,F)$ to a principal bundle, we recall the definition of a principal $G$-bundle. A principal $G$-bundle consists of a total space $P$, a base manifold $B$, a smooth surjective projection $\pi:P\to B$, and a Lie group $G$ acting smoothly, freely, and transitively on the right on each fiber $\pi^{-1}(b)\cong G$, such that $B=P/G$ is the orbit space. The bundle admits local trivializations $\pi^{-1}(U)\cong U\times G$ for an open cover $\{U_{i}\}$ of $B$, with transition functions taking values in $G$.
Theorem 6 (Principal $\mathbb{R}^{n+1}$-bundle structure).
The quantum thermodynamic fiber bundle $(\mathcal{M},\mathcal{B},\\ \Xi,F)$ admits a principal $\mathbb{R}^{n+1}$-bundle structure.
Proof.
Identify the fiber $F\cong\mathbb{R}^{n+1}$ with the additive Lie group $(\mathbb{R}^{n+1},+)$. Define the right action of $G=\mathbb{R}^{n+1}$ on the total space $\mathcal{M}$ by
$(S,\boldsymbol{a},\boldsymbol{\lambda})\cdot g=(S+\delta S,\boldsymbol{a}+\delta\boldsymbol{a},\boldsymbol{\lambda}),$ (70)
for $g=(\delta S,\delta\boldsymbol{a})\in\mathbb{R}^{n+1}$. This action is free, since $p\cdot g=p$ implies $g=(0,\boldsymbol{0})$, the identity. It is transitive on each fiber $F_{\sigma}=\Xi^{-1}(\sigma)$, because for any two points $p_{1}=(S_{1},\boldsymbol{a}_{1},\boldsymbol{\lambda})$, $p_{2}=(S_{2},\boldsymbol{a}_{2},\boldsymbol{\lambda})\in F_{\sigma}$, the group element $g=(S_{2}-S_{1},\boldsymbol{a}_{2}-\boldsymbol{a}_{1})$ satisfies $p_{1}\cdot g=p_{2}$. The action preserves the fibers (fixed $\boldsymbol{\lambda}$) and commutes with the projection $\Xi$, since $\Xi(S,\boldsymbol{a},\boldsymbol{\lambda})=\rho_{\boldsymbol{\mu}(S,\boldsymbol{a},\boldsymbol{\lambda})}$ depends only on $\boldsymbol{\mu}$, which is invariant under translations in $(S,\boldsymbol{a})$.
The base manifold $\mathcal{B}$ is the quotient $\mathcal{M}/\mathbb{R}^{n+1}$, with each orbit  [^1] The quotient map $\mathcal{M}\rightarrow\mathcal{M}/\mathbb{R}^{n+1}=\mathcal{B},\ p\;\mapsto\;[p]=\operatorname{Orb}(p)$ sends every point to its orbit, defined as the set of all points in $\mathcal{M}$ that are reached from a given starting point by acting with the group $\mathbb{R}^{n+1}$. Since the group acts freely and transitively on each fiber, the orbit of any point $p\in\mathcal{M}$ is exactly the whole fiber that contains $p$. (i.e., each equivalence class) corresponding to a fiber $F_{\sigma}$. Local triviality follows from the global coordinate structure $\mathcal{M}\cong\mathbb{R}^{n+1}\times\mathbb{R}^{n}$, where the first factor corresponds to the group $G$ and the second parametrizes $\mathcal{B}$. Thus, $(\mathcal{M},\mathcal{B},\Xi,\mathbb{R}^{n+1})$ is a smooth principal $\mathbb{R}^{n+1}$-bundle with abelian structure group. ∎
This principal structure interprets the redundancy in thermodynamic labeling as a gauge symmetry: group actions correspond to translations in $(S,\boldsymbol{a})$ at fixed $\boldsymbol{\lambda}$, preserving the physical quantum state $\sigma$ but altering its thermodynamic description—analogous to gauge freedom in field theories, where different potentials describe the same physical field.
Definition 22.
A principal connection (or connection $1$-form, or gauge field) on the principal $\mathbb{R}^{n+1}$-bundle $(\mathcal{M},\mathcal{B},\Xi,\mathbb{R}^{n+1})$ is a smooth $\mathfrak{g}$-valued $1$-form $\omega\in\Lambda^{1}(\mathcal{M};\mathfrak{g})$, where $\mathfrak{g}=\mathbb{R}^{n+1}$ is the Lie algebra of the structure group $G=(\mathbb{R}^{n+1},+)$, satisfying the following two properties:
  1. 1.
    For all $g\in G$, $R_{g}^{*}\omega=\mathrm{Ad}_{g^{-1}}\circ\omega$ where $R_{g}:\mathcal{M}\to\mathcal{M}$, $p\mapsto p\cdot g$, is the right action of the group, and $\mathrm{Ad}:G\to\mathrm{Aut}(\mathfrak{g})$ is the adjoint representation. Since $G$ is abelian, $\mathrm{Ad}_{g}Y=Y$ for all $g,Y\in\mathfrak{g}$, so the condition simplifies to
    $R_{g}^{*}\omega=\omega,\ \text{for all }g\in\mathbb{R}^{n+1}.$ (71)
    This means the connection is invariant under global translations in the fiber coordinates $(S,\boldsymbol{a})$.
  2. 2.
    For all $X\in\mathfrak{g}$, let $\xi_{X}\in\mathfrak{X}(\mathcal{M})$ be the fundamental vector field defined by
    $\xi_{X}(p)=\left.\frac{d}{dt}\right|_{t=0}\bigl(p\cdot\exp(tX)\bigr).$ (72)
    Since $\exp(tX)=tX$ for $G=\mathbb{R}^{n+1}$, in coordinates we have $\xi_{X}=X^{0}\partial S+\sum_{i=1}^{n}X^{i}\partial a_{i}$. The connection must satisfy
    $\omega(\xi_{X})=X,\ \text{for all }X\in\mathfrak{g}.$ (73)
    This identifies vertical tangent vectors with infinitesimal group translations.
The principal $\mathbb{R}^{n+1}$-bundle structure $(\mathcal{M},\mathcal{B},\Xi,\mathbb{R}^{n+1})$ reveals that the redundancy in thermodynamic labeling: multiple values of $(S,\boldsymbol{a})$ corresponding to the same physical quantum state $\sigma\in\mathcal{B}$ at fixed intensive parameters $\boldsymbol{\lambda}$ is a gauge symmetry.
Definition 23.
Let $\Xi:\mathcal{M}\to\mathcal{B}$ be the principal quantum thermodynamic bundle. A global gauge transformation is a bundle automorphism $f:\mathcal{M}\to\mathcal{M}$, i.e. a diffeomorphism satisfying:
  1. 1.
    $\Xi\circ f=\Xi$,
  2. 2.
    $f(p\cdot g)=f(p)\cdot g$, for all $p\in\mathcal{M}$, $g\in G$.
The set of all such $f$ forms the gauge group $\mathcal{G}(\mathcal{M})=\mathrm{Aut}(\mathcal{M})$. A local gauge transformation is a bundle automorphism on $\Xi^{-1}(U)\to U$, where $U\subset\mathcal{B}$.
In local coordinates $(S,\boldsymbol{a},\boldsymbol{\lambda})$ on $\mathcal{M}$, a gauge transformation $f\in\mathcal{G}(\mathcal{M})$ acts as
$f(S,\boldsymbol{a},\boldsymbol{\lambda})=\bigl(S+\phi_{S}(\boldsymbol{\lambda}),\boldsymbol{a}+\boldsymbol{\phi}_{\boldsymbol{a}}(\boldsymbol{\lambda}),\boldsymbol{\lambda}\bigr),$ (74)
where $\phi_{S}:\mathcal{B}\to\mathbb{R}$ and $\boldsymbol{\phi}_{\boldsymbol{a}}:\mathcal{B}\to\mathbb{R}^{n}$ are smooth functions. This translates thermodynamic labels within each fiber while preserving the physical state: $\Xi(f(p))=\Xi(p)=\sigma$; and the intensive parameters: $\boldsymbol{\lambda}(f(p))=\boldsymbol{\lambda}(p)$.
Assuming that the Gibbs map is injective, see Remark 9 for details, the following can be stated:
Proposition 12.
The injectivity of the Gibbs map $\boldsymbol{\lambda}\mapsto\rho_{\boldsymbol{\lambda}}$ implies that the equilibrium submanifold $\mathcal{E}$ intersects each fiber $F_{\sigma}$ at exactly one point. Then, equilibrium states are gauge-fixed.
Proof.
For $\sigma=\rho_{\boldsymbol{\lambda}}$, the equilibrium point is $p_{\sigma}=\bigl(S(\sigma),\boldsymbol{a}(\sigma),\boldsymbol{\lambda}\bigr)\in\mathcal{E}$. By injectivity, $\boldsymbol{\lambda}$ is unique. By the maximum entropy principle, $S(\sigma)$ and $\boldsymbol{a}(\sigma)$ are uniquely determined. Any other point in $F_{\sigma}$ has incorrect $(S,\boldsymbol{a})$ and lies off $\mathcal{E}$. ∎
Remark 17.
The horizontal distribution can be defined as the kernel of $\omega$:
$\mathscr{H}_{p}=\ker\omega_{p}=\{v\in T_{p}\mathcal{M}\mid\omega(v)=0\}.$ (75)
This complements the vertical subbundle $\mathscr{V}_{p}=T_{p}F_{\Xi(p)}\cong\mathfrak{g}$ and defines parallel transport.
XI.2 Curvature
In a local trivialization $\Xi^{-1}(U)\cong U\times\mathbb{R}^{n+1}$ of the principal $\mathbb{R}^{n+1}$-bundle $(\mathcal{M},\mathcal{B},\Xi,\mathbb{R}^{n+1})$, let $U\subset\mathcal{B}$ be an open set with coordinates $\{\lambda_{k}\}_{k=1}^{n}$ (so $b_{k}=\lambda_{k}$), and let the fiber $\mathbb{R}^{n+1}$ have coordinates $\{v_{\alpha}\}_{\alpha=0}^{n}$, where $v_{0}=S,$ and $v_{i}=a_{i},\ \text{for }i=1,\dots,n$. The left-invariant Maurer–Cartan form $\theta$ on the abelian Lie group $G=(\mathbb{R}^{n+1},+)$ is
$\theta=\sum_{\alpha=0}^{n}dv_{\alpha}\otimes\dfrac{\partial}{\partial v_{\alpha}}.$ (76)
The principal connection $1$-form $\omega\in\Lambda^{1}(\mathcal{M};\mathfrak{g})$ is defined globally, but it will be useful to consider the local form it takes in our particular trivialization. Concretely, knowing that the horizontal vector fields given by Eq. (60) must satisfy $\omega(e_{k})=0$ for all $k=1,\dots,n$, the connection form can be shown to be given by
$\omega=\sum_{\alpha=0}^{n}\sum_{k=1}^{n}\left(dv_{\alpha}-\Gamma^{k}_{\alpha}d\lambda_{k}\right)\otimes\dfrac{\partial}{\partial v_{\alpha}}.$ (77)
This is the pullback of the Maurer-Cartan form minus the gauge potential $\Gamma^{k}_{\alpha}$. The coefficients $\Gamma^{k}_{\alpha}(\boldsymbol{\lambda})$ encode the geometric coupling between changes in intensive parameters $\boldsymbol{\lambda}$ and shifts in thermodynamic labels $(S,\boldsymbol{a})$ during parallel transport. The connection coefficients $\Gamma^{k}_{\alpha}$ are determined by the orthogonality condition $g_{\mathcal{M}}(e_{k},\partial_{v_{\beta}})=0$. Explicitly:
  • For $\beta=0$, i.e. $\partial_{v_{0}}=\partial_{S}$:
    $g_{\mathcal{M}}\left(e_{k},\dfrac{\partial}{\partial S}\right)=g_{\mathcal{M}}\left(\dfrac{\partial}{\partial\lambda_{k}}-\sum_{\alpha=0}^{n}\Gamma^{k}_{\alpha}\dfrac{\partial}{\partial v_{\alpha}},\dfrac{\partial}{\partial S}\right)=g_{S\lambda_{k}}-\Gamma^{k}_{0}g_{SS}=h_{k}-\Gamma^{k}_{0}g_{S}=0.$ (78)
    Thus, $\Gamma^{k}_{0}=\frac{h_{k}}{g_{S}}$.
  • For $\beta=i\in\{1,\dots,n\}$, i.e., $\partial_{v_{i}}=\partial_{a_{i}}$:
    $\begin{split}g_{\mathcal{M}}\left(e_{k},\dfrac{\partial}{\partial a_{i}}\right)&=g_{\mathcal{M}}\left(\dfrac{\partial}{\partial\lambda_{k}}-\sum_{\alpha=0}^{n}\Gamma^{k}_{\alpha}\dfrac{\partial}{\partial v_{\alpha}},\dfrac{\partial}{\partial a_{i}}\right)=-\Gamma^{k}_{i}g_{a_{i}a_{i}}=-\Gamma^{k}_{i}g_{a_{i}}=0.\end{split}$ (79)
    Thus, $\Gamma^{k}_{i}=0,\text{ for all }i=1,\dots,n$.
Therefore, the connection coefficients are
$\Gamma^{k}_{\alpha}(\boldsymbol{\lambda})=\begin{cases}\displaystyle h_{k}(\boldsymbol{\lambda})/g_{S}(\boldsymbol{\lambda})&\alpha=0\\ 0&\alpha=1,\dots,n\end{cases}$ (80)
Remark 18.
The connection acts exclusively in the entropy direction, reflecting that changes in intensive parameters $\boldsymbol{\lambda}$ induce adjustments in thermodynamic entropy but not in target expectation values, consistent with the structure of Gibbs states.
Proposition 13.
Since $\mathfrak{g}$ is abelian, the curvature $2$-form is $R=d\omega$. In the local trivialization, for horizontal lifts
$e_{k}=\dfrac{\partial}{\partial\lambda_{k}}-\sum_{\alpha=0}^{n}\Gamma^{k}_{\alpha}\dfrac{\partial}{\partial v_{\alpha}},$ (81)
for all $k=1,\ldots,n$, the curvature evaluated on basis vectors is
$R(e_{k},e_{l})=\sum_{\alpha=0}^{n}\left(\dfrac{\partial}{\partial\lambda_{k}}\Gamma^{l}_{\alpha}-\dfrac{\partial}{\partial\lambda_{l}}\Gamma^{k}_{\alpha}\right)\dfrac{\partial}{\partial v_{\alpha}}.$ (82)
Proof.
The local connection form is
$\omega=\sum_{\alpha=0}^{n}\sum_{m=1}^{n}\left(dv_{\alpha}-\Gamma^{m}_{\alpha}d\lambda_{m}\right)\otimes\dfrac{\partial}{\partial v_{\alpha}}.$ (83)
Its exterior derivative is directly given by
$\begin{split}d\omega&=\sum_{\alpha=0}^{n}\sum_{k=1}^{n}\sum_{l=1}^{n}\left(\dfrac{\partial}{\partial\lambda_{k}}\Gamma^{k}_{\alpha}\right)d\lambda_{l}\wedge d\lambda_{k}\otimes\dfrac{\partial}{\partial v_{\alpha}}\\ &=\sum_{\alpha=0}^{n}\sum_{k=1}^{n}\sum_{l=1}^{n}\frac{1}{2}\left(\dfrac{\partial}{\partial\lambda_{k}}\Gamma^{l}_{\alpha}-\dfrac{\partial}{\partial\lambda_{l}}\Gamma^{k}_{\alpha}\right)d\lambda_{k}\wedge d\lambda_{l}\otimes\dfrac{\partial}{\partial v_{\alpha}}.\end{split}$ (84)
Evaluating on the horizontal lifts $e_{k},e_{l}$, the curvature is
$R(e_{k},e_{l})=d\omega(e_{k},e_{l})=\sum_{\alpha=0}^{n}\left(\dfrac{\partial}{\partial\lambda_{k}}\Gamma^{l}_{\alpha}-\dfrac{\partial}{\partial\lambda_{l}}\Gamma^{k}_{\alpha}\right)\dfrac{\partial}{\partial v_{\alpha}},$ (85)
which is Eq. (82). Note that this expression is similar to the general one given by Eq. (62). The Lie bracket vanishes because the group is abelian.∎
Substituting the explicit connection coefficients from Eq. (80), we have that only the $\alpha=0$ term contributes to the curvature:
$R(e_{k},e_{l})$ $=\left[\dfrac{\partial}{\partial\lambda_{k}}\left(\frac{h_{l}}{g_{S}}\right)-\dfrac{\partial}{\partial\lambda_{l}}\left(\frac{h_{k}}{g_{S}}\right)\right]\dfrac{\partial}{\partial S}.$
Corollary 6.
Curvature acts exclusively in the entropy direction, $\partial_{S}$, measuring the non-integrability of the entropy compensation induced by changes in the intensive parameters $\boldsymbol{\lambda}$.
Remark 19.
The connection is flat if and only if
$\dfrac{\partial}{\partial\lambda_{k}}\left(\frac{h_{l}}{g_{S}}\right)=\dfrac{\partial}{\partial\lambda_{l}}\left(\frac{h_{k}}{g_{S}}\right),$ (86)
for all $k,l$, i.e. the $1$-form
$\psi_{i}=\dfrac{h_{i}}{g_{S}}d\lambda_{i},$ (87)
is closed for all $i=1,\ldots,n$.
Corollary 7.
By Proposition 10, the holonomy of a closed loop $\gamma:S^{1}\to\mathcal{B}$ bounding a surface $\mathscr{S}\subset\mathcal{B}$ is
$\mathrm{Hol}(\gamma,p_{0})=-\int_{\mathscr{S}}R,$ (88)
a vertical displacement in the fiber. Non-zero holonomy implies that parallel transport along a cycle in the space of Gibbs states shifts the thermodynamic coordinates $(S,\boldsymbol{a})$, inducing geometric irreversibility $\Sigma_{\gamma}>0$ unless $R=0$.
In this gauge framework, thermodynamic processes are sections of associated vector bundles, and quasistatic transformations correspond to paths with minimal holonomic deviation. Irreversibility in cyclic processes emerges as a curvature effect, unifying thermodynamic dissipation with geometric phases in gauge theories. This mirrors the Aharonov-Bohm phase in electromagnetism or Berry phase in quantum mechanics, where holonomy in a principal bundle over parameter space leads to observable shifts despite local flatness. Here, non-zero $R$ generates non-integrable shifts in $(S,\boldsymbol{a})$, requiring dissipative corrections to close the cycle, manifesting as $\Sigma_{\gamma}>0$—a gauge-induced irreversibility intrinsic to the thermodynamic bundle geometry.
If the principal connection on the quantum thermodynamic bundle is flat, then all cyclic processes in the base manifold $\mathcal{B}$, corresponding to closed loops in intensive parameters $\boldsymbol{\lambda}$, induce zero holonomy. This implies that parallel transport along such cycles returns the system to its initial thermodynamic configuration without any geometric shift in labels $(S,\boldsymbol{a})$, implying no additional entropy production from bundle geometry. In practical terms, flat connections characterize systems where thermodynamic cycles have no gauge-like losses, provided the parameter space $\mathcal{B}$ possesses trivial topology.
Remark 20 (Second law of quantum thermodynamics).
The quantum second law emerges from the non-integrability of the contact structure and the geometry of non-equilibrium paths. For a finite-speed process along a path $\gamma(t)$ in $\mathcal{E}$, the entropy production rate is $\varsigma(t)=\kappa g_{\mathrm{BW}}(\gamma^{\prime}(t),\gamma^{\prime}(t))\geq 0$, with total $\Sigma=\kappa\int_{0}^{T}g_{\mathrm{BW}}(\gamma^{\prime}(t),\gamma^{\prime}(t))dt\geq 0$. Irreversibility is minimal along geodesics and vanishes in the quasistatic limit. Off equilibrium, in fibers $F_{\sigma}$, relaxation paths incur positive entropy production $\Sigma>0$ due to deviations from $\mathcal{E}$. Curvature-induced holonomy in cyclic processes generates additional geometric entropy production.
Remark 21 (Classical thermodynamics as a trivial limit).
Classical thermodynamics is recovered if and only if the fibers $F_{\sigma}$ of the principal bundle $(\mathcal{M},\mathcal{B},\Xi,\mathbb{R}^{n+1})$ are singletons, i.e. $\dim F_{\sigma}=0$ for all $\sigma\in\mathcal{B}$. This trivializes the bundle and makes the group action the identity. Consequently, $\mathcal{M}$ becomes isomorphic to the classical contact manifold $M$. This condition is both necessary and sufficient, as it eliminates quantum redundancies: in quantum systems, the fibers parameterize multiple labels per state, whereas classically, each state has a unique label, causing the fibers to collapse to points. $\Xi$ becomes bijective (the identity on reduced space), aligning the corresponding structures:
  • $\eta$ matches classical Gibbs 1-form,
  • $\mathcal{E}$ reduces to Legendrian $E$,
  • $g_{\mathrm{BW}}$ becomes classical Fisher-Rao.
Mathematically, the bundle trivializes to $\mathcal{M}\cong M\times\{e\}$, where $\{e\}$ is the trivial group. This preserves tangent bundles, contact distributions, and Reeb fields. This criterion unifies the limits, confirming that the quantum theory is the general case and the classical theory emerges as its limit.
XII Conclusion
In this work, we have presented a comprehensive geometrical framework for quantum thermodynamics, grounded in contact geometry and principal fiber bundle theory.
The quantum thermodynamic state space is constructed as a contact manifold, with equilibrium Gibbs states forming Legendrian submanifolds that generalize classical thermodynamic geometry to the quantum regime. The principal fiber bundle structure over the manifold of density operators provides a precise description of non-equilibrium configurations: its fibers characterize relaxation processes, while their unique intersections with the equilibrium submanifold ensure thermodynamic consistency.
The Bures-Wasserstein metric on the equilibrium submanifold provides a Riemannian perspective on quasistatic transformations, minimizing dissipation through geodesics, while diverging geodesic length toward rank-deficient states offers a geometric derivation of the third law, underscoring the unattainability of pure states in finite processes. Extensions to non-equilibrium via pseudo-Riemannian metrics and principal connections reveal curvature-induced holonomy as a source of irreversibility in cyclic processes, quantifying entropy production geometrically.
Quantum thermodynamics is geometry. The laws are not postulates—they are consequences of the manifold, the contact form, and the fiber bundle structure. From the uniqueness of equilibrium to the unattainability of pure states; from reversible geodesics to irreversible holonomy, the full structure of thermal behavior in quantum systems is encoded in the differential and fiber geometry of the quantum thermodynamic state space.
XIII Acknowledgements
The authors wish to express their gratitude to Prof. Frederic Schuller, whose lectures inspired them to pursue studies in theoretical and mathematical physics during their undergraduate years.
We acknowledge funding from the Ministry for Digital Transformation and of Civil Service of the Spanish Government through projects PID2021-126217NB-I00, PID2023-149365NB-I00, PID2024-162155OB-I00, FPU20/02835 and QUANTUM ENIA project call - Quantum Spain project, and by the European Union through the Recovery, Transformation and Resilience Plan - NextGenerationEU within the framework of the Digital Spain 2026 Agenda. Finally, we are also grateful for the technical support provided by PROTEUS, the supercomputing center of the Institute Carlos I for Theoretical and Computational Physics in Granada, Spain.